### v2

# Feature learning in finite-width Bayesian deep linear networks with multiple outputs and convolutional layers

Abstract
Deep linear networks have been extensively studied, as they provide simplified models of deep learning. However, little is known in the case of finite-width architectures with multiple outputs and convolutional layers. In this manuscript, we provide rigorous results for the statistics of functions implemented by the aforementioned class of networks, thus moving closer to a complete characterization of feature learning in the Bayesian setting. Our results include: (i) an exact and elementary non-asymptotic integral representation for the joint prior distribution over the outputs, given in terms of a mixture of Gaussians; (ii) an analytical formula for the posterior distribution in the case of squared error loss function (Gaussian likelihood); (iii) a quantitative description of the feature learning infinite-width regime, using large deviation theory. From a physical perspective, deep architectures with multiple outputs or convolutional layers represent different manifestations of kernel shape renormalization, and our work provides a dictionary that translates this physics intuition and terminology into rigorous Bayesian statistics.
Keywords: Deep learning theory, Bayesian deep linear networks, Convolutional layers, Feature learning, Gaussian mixtures
1 Introduction
Deep learning makes use of a wide array of architectures, each with unique strengths and applications. Selecting the optimal architecture for a specific task requires a deep understanding of their properties and a principled way to compute them from theory. Unfortunately, the empirical advancements in this field risk outpacing the development of a comprehensive theoretical framework.
Significant progress has been made in certain asymptotic regimes, where neural networks enjoy forms of universality that simplify their analysis. In the (lazy-training) infinite-width limit, in particular, universality is manifested in both the training under gradient flow, described by the neural tangent kernel (Jacot et al., 2018), and the Bayesian inference setting (Lee et al., 2018b), where exact relations between neural networks and kernel methods were obtained. However, these results cannot explain the success of modern deep architectures, which perform non-trivial feature selection beyond what is possible in the lazy-training regime (Chizat et al., 2019; Lewkowycz et al., 2021; Brown et al., 2020). One can circumvent this problem by considering different rescalings of the weights with the width of the layers, using the so-called mean field (Mei et al., 2018; Rotskoff and Vanden-Eijnden, 2022; Sirignano and Spiliopoulos, 2020; Chizat and Bach, 2018) or the more recent maximal update parametrizations (Yang and Hu, 2021).
Despite these insights, neural networks at finite width, which are relevant for applications, display complex and intriguing phenomena that are not captured by infinite-width asymptotics (Ciceri et al., 2024; Ingrosso and Goldt, 2022; Petrini et al., 2023b, a; Sclocchi and Wyart, 2024; Sclocchi et al., 2023). Addressing these finite-size properties calls for a non-perturbative theory, one which does not rely on large-size limits. Developing such a theory for general neural networks has proven to be quite challenging. However, progress has been made by considering simplified models where the activation functions are linear. Deep linear networks (Saxe et al., 2014, 2019; Li and Sompolinsky, 2021) are linear in terms of their inputs but they retain non-linear dependence on the parameters. They provide analytically tractable non-convex problems in parameter space, thus serving as a bridge between idealized infinite-width models and the reality of finite-sized networks.
Here, we address the computation of network statistics in the Bayesian learning setting, equivalent to the statistical mechanics framework employed by physicists (Neal, 1996; Lee et al., 2018a; Novak et al., 2019; Garriga-Alonso et al., 2019; Engel and Van den Broeck, 2001). In Bayesian deep learning, the prior over the network’s parameters induces a prior over the network’s outputs (more details are in Sec. 2.3). In the (lazy-training) infinite-width limit, and for non-linear activation function, the prior over the outputs is known to be Gaussian both for fully connected architectures (Neal, 1996; de G. Matthews et al., 2018; Hanin, 2023) and for convolutional architectures (Novak et al., 2019; Garriga-Alonso et al., 2019). For fully connected networks, precise rates of convergence to normality have been investigated as well (see Favaro et al., 2023; Trevisan, 2023). The goal of our work is to characterize the non-Gaussian behavior, at finite depth and width, of two classes of deep linear networks: (i) those with fully-connected layers and multiple outputs, and (ii) those with convolutional layers and a single linear readout. Our results, summarized in the next section, provide a way to quantify non-perturbative feature learning effects in these architectures. We will discuss the link between our work and the literature by physicists in Sec. 4.
1.1 Informal statement of the results
We summarize here our main results informally, as three take-home messages.
Take-home message 1: At finite width, the output prior is an exactly computable mixture of Gaussians. The sizes of the hidden layers appear as parameters in the mixing measure, leading to dimensional reduction.
We compute the prior over the outputs at finite-width in the linear case, showing that it is a mixture of Gaussians with an explicit mixing distribution. Notably, in this representation the covariance, i.e. the neural network Gaussian process (NNGP) kernel, is modified by $L$ (the number of hidden layers) random matrices with Wishart distribution. For fully-connected linear networks with finite number $D$ of outputs in the readout layer, the dimension of these matrices is $D$ (Proposition 1), while for convolutional linear networks (with unitary stride), the dimension of the Wishart matrices is the (fixed) size of the input $N_{0}$ (Proposition 5). All the sizes of the hidden layers $N_{\ell}$ appear parametrically in the prior, providing an explicit dimensional reduction.
Take-home message 2: At finite width the posterior predictive is a mixture of Gaussians with closed form mixing distribution.
In the case of quadratic loss function (Gaussian likelihood), the posterior distribution inherits the properties of the prior. The posterior is again a mixture of Gaussians and this leads to the rather standard equations for the bias and variance of a Gaussian Process (Rasmussen and Williams, 2005), with the important difference that they are now random variables (Proposition 6 and 7).
Take-home message 3: In the feature learning infinite-width limit, large deviation asymptotics shows non-trivial explicit dependence on the training inputs and labels.
Using the simple parametric dependence of our formulas from the layer widths, we provide an asymptotic analysis in the limit of large width, using the language of Large Deviation Theory. First, we recover the well-known infinite-width limit, showing that the prior is degenerate in this case, since the Wishart ensembles concentrate around the identity matrix (Proposition 8). Second, we consider the so-called feature learning infinite-width limit (Chizat et al., 2019; Mei et al., 2018; Geiger et al., 2021, 2020; Yang and Hu, 2021), which is equivalent to a different re-parametrization of the loss and of the output function, and we precisely show how it provides a way to escape lazy training. The measure over the Wishart ensembles concentrates also in this case, but this time they do so around non-trivial solutions that explicitly depend on the training inputs and labels (Proposition 10).
1.2 Related work
Saxe et al. (2014, 2019) derived exact solutions for the training dynamics of deep linear networks (DLN). More recently, Li and Sompolinsky (2021) studied deep linear networks in a Bayesian framework, via the equivalent statistical mechanics formulation, and obtained analytical results in the proportional limit, where the common size of the hidden layers $N$ and the size of the training set $P$ are taken to infinity, while keeping their ratio $\alpha=P/N$ fixed. This analysis was later generalized to globally gated DLNs (Li and Sompolinsky, 2022), and extended to the data-averaged case, using the replica method from spin glass theory (Zavatone-Veth et al., 2022).
Hanin and Zlokapa (2023) reconsidered the same setting in the case of single-output networks, and they derived a non-asymptotic result for the Bayesian model evidence (partition function in statistical mechanics), in terms of Meijer G-functions (the fact that the prior over a single output is related to these special functions was shown in Zavatone-Veth and Pehlevan, 2021). They employed this result to investigate various asymptotic limits where $N$, $P$, and the depth of the network $L$ are simultaneously taken to infinity at the same rate.
Closely related to the deep linear literature, Pacelli et al. (2023) proposed an effective theory for deep non-linear networks in the proportional limit, whose derivation is based on a Gaussian equivalence (Goldt et al., 2020, 2022; Gerace et al., 2020; Hu and Lu, 2023; Gerace et al., 2024; Aguirre-López et al., 2024) heuristically justified using Breuer-Major theorems (Breuer and Major, 1983; Bardet and Surgailis, 2013; Nourdin et al., 2011). This effective theory has been tested with great accuracy for one-hidden-layer fully-connected networks with a single output in Baglioni et al., 2024. Camilli et al., 2023 and Cui et al., 2023 reconsidered the same setting in the Bayes optimal and data-averaged case. Other non-perturbative approaches to Bayesian deep non-linear networks include Naveh and Ringel (2021); Seroussi et al. (2023); Fischer et al. (2024) and the recent renormalization group framework proposed in Howard et al. (2024a, b).
1.3 Kernel renormalization in the proportional limit
Bayesian networks in the infinite-width limit ($N\to\infty$ at finite $P$) are equivalent to Gaussian processes (NNGP), whose covariance is given by a kernel (Lee et al., 2018a). A powerful view of Bayesian networks in the proportional limit (Li and Sompolinsky, 2021; Pacelli et al., 2023; Li and Sompolinsky, 2022) focuses on how the prior is modified layer by layer. For fully-connected architectures with a single output, the NNGP kernel at each layer is rescaled by a real number (a scalar order parameter in statistical mechanics). The order parameters satisfy a set of coupled equations that explicitly depend on the inputs and labels in the training set. This process is referred to as a scalar renormalization of the kernel. In fully-connected networks with multiple outputs, instead, the order parameters are $D\times D$ matrices ($D$ being the size of the readout layer), and they modify the NNGP kernel via a Kronecker product. This process is called kernel shape renormalization (Li and Sompolinsky, 2021; Pacelli et al., 2023). Yet another transformation of the kernel, coined local kernel renormalization, was found for convolutional networks in Aiudi et al. (2023), where the NNGP kernel splits into several local components (Novak et al., 2019), labelled by two additional indices $i,j$ that run across the patches in each element of the training set (e.g., the receptive fields of an image). In the proportional limit, each local component $(i,j)$ of the kernel is modified by an order parameter $Q_{ij}$. An interesting type of kernel shape renormalization was found in globally gated deep linear networks too (Li and Sompolinsky, 2022), but the interpretation is different than in the convolutional case presented in Aiudi et al. (2023).
Common definitions
$P$ size of the training set
$L$ depth of the network
$N_{0}$ input dimension
$\theta$ collection of all trainable weights, $\theta=\{W^{(\ell)}\}_{\ell=0}^{L}$
$\mathbf{x}^{\mu}$, $X$ input of the network in the training set, $X=[\mathbf{x}^{\mu}]_{\mu=1}^{P}$
$\mathbf{y}^{\mu}$, $\mathbf{y}_{1:P}$ label (response) of $\mathbf{x}^{\mu}$ in the training set, $\mathbf{y}_{1:P}^{\top}=(\mathbf{y}^{1\top},\dots,\mathbf{y}^{P\top})$
$f_{\theta}(\mathbf{x}^{\mu})$ output of the network from input $\mathbf{x}^{\mu}$
Fully-connected
$N_{1},\cdots,N_{L}$ width of each hidden layer
$N_{L+1}=D$ output dimension
$\mathbf{S}^{\mu}$, $\mathbf{S}_{1:P}$ output of the network in $\mathbb{R}^{D}$, $\mathbf{S}^{\mu}=f_{\theta}(\mathbf{x}^{\mu})$, $\mathbf{S}_{1:P}^{\top}=(\mathbf{S}^{1\top},\dots,\mathbf{S}^{P\top})$
Convolutional
$C_{0},\cdots,C_{L}$ number of channels in each hidden layer
$M$ dimension of the channel mask
$S^{\mu}$, $\mathbf{S}_{1:P}$ output of the network in $\mathbb{R}$, $S^{\mu}=f_{\theta}(\mathbf{x}^{\mu})$, $\mathbf{S}_{1:P}^{\top}=(S^{1},\dots,S^{P})$
Table 1: Table of notations
2 Problem setting
We consider a supervised learning problem with training set $\{\mathbf{x}^{\mu},\mathbf{y}^{\mu}\}_{\mu=1}^{P}$, where each $\mathbf{x}^{\mu}\in\mathbb{R}^{N_{0}\times C_{0}}$ and the corresponding labels (response) $\mathbf{y}^{\mu}\in\mathbb{R}^{D}$, with $C_{0}=1$ for the fully-connected architecture and, for simplicity, $D=1$ for the convolutional architecture. The output will be denoted by $f_{\theta}(\mathbf{x}^{\mu})$, where $\theta$ represents the collection of all the trainable weights of the network (see Table 1 for a summary of notations).
2.1 Fully-connected deep linear neural networks (FC-DLNs)
In fully-connected neural networks, the pre-activations of each layer $h_{i_{\ell}}^{(\ell)}$ ($i_{\ell}=1,\dots,N_{\ell}$; $\ell=1,\dots,L$) are given recursively as a function of the pre-activations of the previous layer $h_{i_{\ell-1}}^{(\ell-1)}$ ($i_{\ell-1}=1,\dots,N_{\ell-1}$):
$\begin{split}h_{i_{1}}^{(1)}(\mathbf{x}^{\mu})&=\frac{1}{\sqrt{N_{0}}}\sum_{i_{0}=1}^{N_{0}}W^{(0)}_{i_{1}i_{0}}x_{i_{0}}^{\mu}+b_{i_{1}}^{(0)}\,,\\ h_{i_{\ell}}^{(\ell)}(\mathbf{x}^{\mu})&=\frac{1}{\sqrt{N_{\ell-1}}}\sum_{i_{\ell-1}=1}^{N_{\ell-1}}W^{(\ell-1)}_{i_{\ell}i_{\ell-1}}h_{i_{\ell-1}}^{(\ell-1)}(\mathbf{x}^{\mu})+b_{i_{\ell}}^{(\ell-1)}\,,\\ \end{split}$ (1)
where $W^{(\ell-1)}$ and $b^{(\ell-1)}$ are respectively the weights and the biases of the $\ell$-th layer, whereas the input layer has dimension $N_{0}$ (the input data dimension). In what follows we shall take $b_{i_{\ell}}^{(\ell-1)}=0$ for every $\ell$ and every $i_{\ell}$ (without loss of generality, as one can include this contribution in the vectors of weights adding additional dimensions in input and hidden space).
We add one last readout layer with dimension $N_{L+1}=D$, and define the function implemented by the deep neural network as $f_{\theta}(\mathbf{x}^{\mu})=(S^{\mu}_{1},\dots,S^{\mu}_{D})^{\top}$, where
$S^{\mu}_{i}=\frac{1}{\sqrt{N_{L}}}\sum_{i_{L}=1}^{N_{L}}W^{(L)}_{i,i_{L}}h_{i_{L}}^{(L)}(\mathbf{x}^{\mu})\qquad i=1,\dots,D\,;\,\,\mu=1,\dots,P,$ (2)
$W^{(L)}_{i,i_{L}}$ being the weights of the last layer. The input training set is collected in a $N_{0}\times P$ matrix $X$, that is
$X=[\mathbf{x}^{1},\dots,\mathbf{x}^{P}],$
and the corresponding labels in a vector $\mathbf{y}_{1:P}^{\top}=(\mathbf{y}_{1}^{\top},\dots,\mathbf{y}_{P}^{\top})$. We also denote by
$\mathbf{S}_{1:P}=(S^{1}_{1},\dots,S^{1}_{D},S^{2}_{1},\dots,S^{P}_{1},\dots,S^{P}_{D})^{\top}$
the $P$ outputs $(\mathbf{S}^{\mu}=f_{\theta}(\mathbf{x}^{\mu}):\mu=1,\dots,P)$ stacked in a vector.
2.2 Convolutional deep linear neural networks (C-DLNs)
For convolutional neural networks, the pre-activations at each layer $\ell=1,\dots,L$ are labelled by two indices: (i) the channel index $a_{\ell}=1,\dots,C_{\ell}$, where $C_{\ell}$ is the total number of channels in each layer; (ii) the spatial index $i=1,\dots,N_{0}$, which runs over the input coordinates. Similarly any input has two indices, that is $\mathbf{x}^{\mu}=[x^{\mu}_{a_{0},i}]$, with $a_{0}=1,\dots,C_{0}$ and $i-1,\dots,N_{0}$. Pre-activations of the first layer are given by
$h_{a_{1},i}^{(1)}(\mathbf{x}^{\mu})=\frac{1}{\sqrt{MC_{0}}}\sum_{a_{0}=1}^{C_{0}}\sum_{m=-\lfloor M/2\rfloor}^{\lfloor M/2\rfloor}W^{(0)}_{m,a_{1}a_{0}}x^{\mu}_{a_{0},i+m(\mathrm{mod}N_{0})}$ (3)
and for $\ell>1$
$h_{a_{\ell},i}^{(\ell)}(\mathbf{x}^{\mu})=\frac{1}{\sqrt{MC_{\ell-1}}}\sum_{a_{\ell-1}=1}^{C_{\ell-1}}\sum_{m=-\lfloor M/2\rfloor}^{\lfloor M/2\rfloor}W^{(\ell-1)}_{m,a_{\ell}a_{\ell-1}}h^{(\ell-1)}_{a_{\ell-1},i+m(\mathrm{mod}N_{0})}(\mathbf{x}^{\mu})\,,$ (4)
where $M$ is the dimension of the channel mask. Here we restrict our analysis to $1d$ convolutions, periodic boundary conditions over the spatial index and unitary stride. This choice allows to keep the size of the spatial index constant through layers. Note that in the case of non-unitary stride, one should introduce a different spatial index $i_{\ell}$ for each layer, which runs from 1 to the integer part of the ratio between $N_{0}$ and the stride’s $\ell$th power. The output of the network is given by
$S^{\mu}=f_{\theta}(\mathbf{x}^{\mu})=\frac{1}{\sqrt{C_{L}N_{0}}}\sum_{a_{L}=1}^{C_{L}}\sum_{i=1}^{N_{0}}W^{(L)}_{a_{L}i}h^{(L)}_{a_{L},i}(\mathbf{x}^{\mu})\,.$ (5)
Again we denote by $\mathbf{S}_{1:P}$ the vector of outputs $(S^{1},\dots,S^{P})$ and the data is now collected in a three-dimensional array $X=[x^{\mu}_{a_{0},i}]_{\mu,a_{0},i}$.
2.3 Bayesian setting
In a Bayesian neural network, a prior for the weights $\theta$ is specified, which translates in a prior for $f_{\theta}(\mathbf{x}^{\mu})$. In other words, the prior (over the outputs) is the density of the random vector $\mathbf{S}_{1:P}$. This density evaluated for a given value $\mathbf{s}_{1:P}$ taken by $\mathbf{S}_{1:P}$, will be denoted by $p_{\mathrm{prior}}(\mathbf{s}_{1:P}|X)$. At this stage, one needs also a likelihood for the labels given the inputs and the outputs, denoted by $\mathcal{L}(\mathbf{y}_{1:P}|\mathbf{s}_{1:P},\beta)$, where $\beta$ is an additional parameter, which plays the role of the inverse temperature in statistical physics. The posterior density of $\mathbf{S}_{1:P}$ given $\mathbf{y}_{1:P}$ is (by Bayes theorem)
$p_{\mathrm{post}}(\mathbf{s}_{1:P}|\mathbf{y}_{1:P},X)\propto p_{\mathrm{prior}}(\mathbf{s}_{1:P}|X)\mathcal{L}(\mathbf{y}_{1:P}|\mathbf{s}_{1:P}).$
In order to describe the posterior predictive, let $\mathbf{x}^{0}$ be a new input and set $\mathbf{S}^{0}=f_{\theta}(\mathbf{x}^{0})$. The posterior predictive density of $\mathbf{S}_{0}$ given $\mathbf{y}_{1:P}$ is obtained similarly by
[…]
where $p_{\mathrm{prior}}(\mathbf{s}_{0},\mathbf{s}_{1:P}|\mathbf{x}^{0},X)$ is the prior for $(\mathbf{S}_{0},\mathbf{S}_{1:P})$ and
[…]
is the partition function (normalization constant). In the convolutional setting we have $D=1$ and, consequently, we shall write $s_{0}$ in place of $\mathbf{s}_{0}$.
Note that, training the network with a quadratic loss function is equivalent to consider a Gaussian likelihood in the Bayesian setting, that is
$\mathcal{L}(\mathbf{y}_{1:P}|\mathbf{s}_{1:P},\beta)\propto e^{-\frac{\beta}{2}\|\mathbf{s}_{1:P}-\mathbf{y}_{1:P}\|^{2}}.$
Note that in this case
$\mathbf{y}^{\mu}=\mathbf{S}^{\mu}+\bm{\epsilon}^{\mu}=f_{\theta}(\mathbf{x}^{\mu})+\bm{\epsilon}^{\mu}\quad\bm{\epsilon}^{\mu}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\beta^{-1}\mathbb{1}_{D}).$
Here, we assume that the weights are independent normally distributed with zero means and variance $1/\lambda_{\ell}$ at layer $\ell$. In summary:
  • For the fully connected network (FC-DLN):
    $W^{(\ell)}_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\lambda_{\ell}^{-1})$ (6)
  • For the convolutional network (C-DLN):
    $W^{(\ell)}_{m,ab}\overset{iid}{\sim}\mathcal{N}(0,\lambda_{\ell}^{-1})\,,\qquad W^{(L)}_{ai}\overset{iid}{\sim}\mathcal{N}(0,\lambda_{L}^{-1})\,.$ (7)
3 Results
Our main results are collected in this section.
3.1 Exact non-asymptotic integral representation for the prior over the outputs
We start discussing the FC-DLN. To state the results, denote by $\mathcal{S}^{+}_{D}$ the set of symmetric positive definite matrices on $\mathbb{R}^{D}$ and recall that a random matrix $Q$ taking values in $\mathcal{S}^{+}_{D}$ has Wishart distribution with $N>D$ degrees of freedom and scale matrix $V$ if it has the following density (with respect to the Lebesgue measure on the cone of symmetric positive definite matrices)
$\mathcal{W}_{D}(Q|V,N)=\frac{\det(Q)^{\frac{N-D-1}{2}}e^{-\frac{1}{2}\operatorname{tr}(V^{-1}Q)}}{\det(V)^{N/2}2^{DN/2}(\pi)^{D(D-1)/4}\prod_{k=1}^{D}\Gamma\!\left(\frac{N-D+k}{2}\right)}.$
Equivalently, $Q$ has a Wishart distribution with $N$ degrees of freedom and scale matrix $V$ if its law corresponds to the law of $\sum_{i=1}^{N}\mathbf{Z}_{i}\mathbf{Z}_{i}^{\top}$ where $\mathbf{Z}_{i}$ are independent Gaussian vectors in $\mathbb{R}^{D}$ with zero mean and covariance matrix $V$ (see Eaton, 2007 and the references therein). If $D=1$, one has that $\mathcal{S}^{+}_{D}=\mathbb{R}^{+}$ and the Wishart distribution reduces to a Gamma distribution of parameters $\alpha=\frac{N}{2}$, $\beta=\frac{1}{2V}$, that is
$\mathcal{W}_{1}(Q|V,N)=\frac{1}{(2V)^{N/2}\Gamma(N/2)}Q^{\frac{N}{2}-1}e^{-\frac{1}{2V}Q}\quad\text{for $Q>0$}.$
In our integral representations a fundamental role is played by the joint distribution of the vector of matrices
$(Q_{1},\dots,Q_{L}):=(\tilde{Q}_{1}/N_{1},\dots,\tilde{Q}_{L}/N_{L}),$
where the $\tilde{Q}_{\ell}$’s are $D\times D$ independent Wishart random matrices with $N_{\ell}$ degree of freedoms and identity scale matrix $\mathbb{1}_{D}$. This joint distribution of $(Q_{1},\dots,Q_{L})$ is
$\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L}):=\left(\prod_{\ell=1}^{L}\mathcal{W}_{D}\Big{(}Q_{\ell}\Big{|}\frac{1}{N_{\ell}}\mathbb{1}_{D},N_{\ell}\Big{)}\right)dQ_{1}\dots dQ_{L}$ (8)
For the sake of notation, we write $\mathcal{Q}_{L,N}$ in place of the more correct $\mathcal{Q}_{L,N_{1},\dots,N_{L}}$.
We now show that the prior distribution over $\mathbf{S}_{1:P}$ is a mixture of Gaussians where the covariance matrix is an explicit function of random Wishart matrices. In what follows $A\otimes B$ denotes the Kronecker product of matrices $A$ and $B$. Moreover, we set $\lambda^{*}:=\lambda_{0}\dots\lambda_{L}$.
Proposition 1
(non-asymptotic integral representation for the prior of FC-DLNs with multiple outputs) Let $\hat{p}_{\mathrm{prior}}$ be the characteristic function of the outputs $\mathbf{S}_{1:P}$ of a FC-DLN, that is $\hat{p}_{\mathrm{prior}}(\bar{\mathbf{s}}_{1:P}|X)=\mathbf{E}[\exp\{i\bar{\mathbf{s}}_{1:P}^{\top}\mathbf{S}_{1:P}\}]$. If $\min(N_{\ell}:\ell=1,\dots,L)>D$, then
$\hat{p}_{\mathrm{prior}}(\bar{\mathbf{s}}_{1:P}|X)=\int_{(\mathcal{S}^{+}_{D})^{L}}\!\!\!e^{-\frac{1}{2}\bar{\mathbf{s}}_{1:P}^{\top}\mathcal{K}_{\mathrm{FC}}(Q_{1},\dots,Q_{L})\bar{\mathbf{s}}_{1:P}}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L})$
where $\mathcal{Q}_{L,N}$ is defined by (8), $\mathcal{K}_{\mathrm{FC}}(Q_{1},\dots,Q_{L})=(N_{0}\lambda^{*})^{-1}X^{\top}X\otimes Q^{(L)}(Q_{1},\dots,Q_{L})$, and
$Q^{(L)}=Q^{(L)}(Q_{1},\dots,Q_{L}):=(U_{1}\dots U_{L})^{\top}U_{1}\dots U_{L},$
$U_{\ell}$ being $D\times D$ matrices such that $U_{\ell}^{\top}U_{\ell}=Q_{\ell}$.
Remark 2
In the previous representation one can choose any decomposition of $Q_{\ell}$ in $U_{\ell}^{\top}U_{\ell}$. For convenience in what follows we choose the Cholesky decomposition where $U_{\ell}^{\top}=U_{\ell}^{\top}(Q_{\ell})$ is a lower triangular matrix with positive diagonal entries. This $U_{\ell}^{\top}$ is called Cholesky factor (or Cholesky square root). The Cholesky square root is one to one and continuous from $\mathcal{S}_{D}^{+}$ to its image. When $Q_{\ell}$ has a Wishart distribution the law of it Cholesky square root $U_{\ell}^{\top}$ is the so-called Bartlett distribution (see Kshirsagar, 1959).
Remark 3
The integral representation stated above is equivalent to say that the random matrix $S=[S^{\mu}_{d}]_{d=1,\dots,D;\mu=1,\dots,P}$ admits the explicit stochastic representation
$S=U_{L}^{\top}\cdots U_{1}^{\top}\frac{Z}{\sqrt{N_{0}\lambda^{*}}}X$ (9)
where $U_{\ell}$ are $D\times D$ independent random matrices such that $Q_{\ell}=U_{\ell}^{\top}U_{\ell}$ has a Wishart distribution with $N_{\ell}$ degrees of freedom and scale matrix $\frac{1}{N_{\ell}}\mathbb{1}_{D}$ and $Z$ is a $D\times N_{0}$ matrix of independent standard normal random variables. As mentioned in the introduction, all the $N_{\ell}$’s appear parametrically in the above representation, providing the anticipated dimensional reduction. For comparison, note that by (1) one can derive another explicit explicit expression for $S$, namely
$S=\frac{W^{(L)}}{\sqrt{N_{L}}}\cdots\frac{W^{(0)}}{\sqrt{N_{0}}}X$ (10)
where $W^{(\ell)}=[W_{i_{\ell+1},i_{\ell}}^{(\ell)}]$ are $N_{\ell+1}\times N_{\ell}$ matrices of normal random variables of zero mean and variance $1/\lambda_{\ell}$. This second representation is more direct (and easier to derive) and also shows that $S$ is a mixture of Gaussians. Nevertheless, containing the $N_{\ell+1}\times N_{\ell}$ matrices $W^{(\ell)}$, it does not provide any dimensional reduction.
Remark 4
As recalled above, for $D=1$ the $U_{\ell}$’s are independent random variables distributed as the square root of a $\operatorname{Gamma}(N_{\ell}/2,N_{\ell}/2)$ random variable, that is $N_{\ell}{U_{\ell}}^{2}$ is a $\chi^{2}$ with ${N_{\ell}}$ degrees of freedom. In the case of Gaussian likelihood, we can employ this result to obtain an alternative integral representation of the non-asymptotic partition function (Bayesian model evidence) given by Hanin and Zlokapa (2023).
While in Hanin and Zlokapa (2023) the partition function is given in terms of Meijer G-functions, here it is expressed as an $L-$dimensional integral over a set of order parameters $Q_{\ell}$. Using our alternative representation, it is possible to compare and find an equivalence with the results (in the proportional limit $P,N_{\ell}\to\infty$ at fixed $\alpha_{\ell}=P/N_{\ell}$) found in Li and Sompolinsky (2021) and Pacelli et al. (2023) (see also Discussion in Section 4). Note that the link between Meijer G-functions and the integral representation based on Gamma functions was first established in Zavatone-Veth and Pehlevan (2021) in the limited setting of individual training patterns (i.e. for the prior over a single input vector). It should be noted that when $L=1$, Proposition 1 shows that $\mathbf{S}_{1:P}$ is a mixture of Gaussians with a Wishart matrix as covariance. It is known that the resulting distribution is the so-called multivariate Student’s $t$-distribution (see, e.g., Representation B in Lin, 1972). One can show that the characteristic function of a multivariate Student’s $t$ can be expressed in term of Macdonald functions (Joarder and Alam, 1995). Since the Macdonald function is recovered as a special case of a Meijer G-function, this clarify (at least for $L=1$) the link between our stochastic representation and the Fourier representation given in Zavatone-Veth and Pehlevan (2021) for $P=1$.
In order to discuss a non-asymptotic representation formula for the convolutional architecture, it is convenient to introduce the translation operator $T_{m}$ as
$T_{m,ij}=\delta_{j,i+m(\mathrm{mod}N_{0})}.$ (11)
In this way, Eq. (4) takes the following form
$h_{a_{\ell},i}^{(\ell)}(\mathbf{x}^{\mu})=\frac{1}{\sqrt{MC_{\ell-1}}}\sum_{a_{\ell-1}=1}^{C_{\ell-1}}\sum_{m=-\lfloor M/2\rfloor}^{\lfloor M/2\rfloor}W^{(\ell-1)}_{m,a_{\ell}a_{\ell-1}}\sum_{j=1}^{N_{0}}T_{m,ij}h^{(\ell-1)}_{a_{\ell-1},j}(\mathbf{x}^{\mu})\,.$ (12)
One can now define a map $\mathcal{T}:(\mathcal{S}^{+}_{N_{0}})^{L}\to\mathcal{S}^{+}_{N_{0}}$ that will play a role similar to the map $Q^{(L)}=Q^{(L)}(Q_{1},\dots,Q_{L})$ introduced in Proposition 1. Given $Q_{1},\dots,Q_{L}$, one sets $\mathcal{T}(Q_{1},\dots,Q_{L})=Q_{1}^{*}$ where $Q_{L}^{*},\dots,Q_{1}^{*}$ are defined by the following backward recursion:
$Q^{*}_{L}$ $=\frac{1}{M}\sum_{m=-\lfloor M/2\rfloor}^{\lfloor M/2\rfloor}T^{\top}_{m}Q_{L}T_{m}\,,$ (13)
${Q}_{\ell}^{*}$ $=\frac{1}{M}\sum_{m=-\lfloor M/2\rfloor}^{\lfloor M/2\rfloor}T^{\top}_{m}\left((U_{\ell+1}^{*})^{\top}Q_{\ell}U_{\ell+1}^{*}\right)T_{m}\quad\text{for $\ell=L-1,\dots,1$}$
$U_{\ell+1}^{*}$ being a square root of $Q^{*}_{\ell+1}$, that is $(U_{\ell+1}^{*})^{\top}U_{\ell+1}^{*}=Q_{\ell+1}^{*}$. Note that $[T_{m}^{\top}QT_{m}]_{i,j}=Q_{i+m(\text{mod}N_{0}),j+m(\text{mod}N_{0})}$ for any matrix $Q$. Finally, we need the analogous of $\mathcal{Q}_{L,N}$ for the convolutional architecture. To this end, we set
$\mathcal{Q}_{L,C}(dQ_{1}\dots dQ_{L}):=\left(\prod_{\ell=1}^{L}\mathcal{W}_{N_{0}}\Big{(}Q_{\ell}\Big{|}\frac{1}{C_{\ell}}\mathbb{1}_{N_{0}},C_{\ell}\Big{)}\right)dQ_{1}\dots dQ_{L}.$ (14)
Note that $\mathcal{Q}_{L,C}$ has exactly the same form of $\mathcal{Q}_{L,N}$ once $N_{1},\dots,N_{L}$ is replaced by $C_{1},\dots,C_{L}$.
Proposition 5
(non-asymptotic integral representation for the prior of C-DLNs with single output) For a C-DLN with $\min(C_{\ell}:\ell=1,\dots,L)>N_{0}$, the characteristic function $\hat{p}_{\mathrm{prior}}(\bar{\mathbf{s}}_{1:P}|X)=\mathbf{E}[\exp\{i\bar{\mathbf{s}}_{1:P}^{\top}\mathbf{S}_{1:P}\}]$ of the outputs $\mathbf{S}_{1:P}$ is given by
$\hat{p}_{\mathrm{prior}}(\bar{\mathbf{s}}_{1:P}|X)=\int_{(\mathcal{S}^{+}_{N_{0}})^{L}}e^{-\frac{1}{2}\bar{\mathbf{s}}_{1:P}^{\top}(\mathcal{K}_{\mathrm{C}}(Q_{1},\dots,Q_{L}))\bar{\mathbf{s}}_{1:P}}\mathcal{Q}_{L,C}(dQ_{1}\dots dQ_{L})$
where $\mathcal{Q}_{L,N}$ is defined by (14), and the covariance matrix $\mathcal{K}_{\mathrm{C}}(Q_{1},\dots,Q_{L})$ is defined as
$\mathcal{K}_{\mathrm{C}}(Q_{1},\dots,Q_{L}):=\Big{[}\sum_{r,s}\mathcal{T}(Q_{1},\dots,Q_{L})_{rs}\sum_{a_{0}}\frac{x^{\mu}_{a_{0},r}x^{\nu}_{a_{0},s}}{\lambda^{*}C_{0}}\Big{]}_{\mu=1,\dots,P;\nu=1,\dots,P}.$ (15)
Note that the map $\mathcal{T}$ acts as on a four indices tensor, providing a partial trace wrt to indices $r,\,s$. This tensor is known in the literature as local covariance matrix from the seminal work Novak et al. (2019) on the Bayesian infinite-width limit in convolutional networks. The same partial trace structure has been also identified in the proportional limit for non-linear one-hidden-layer networks in Aiudi et al. (2023), where the authors coined the term local kernel renormalization to indicate the re-weighting of the infinite-width local NNGP kernel found in Novak et al. (2019).
3.2 Posterior distribution and predictor’s statistics in the case of squared error loss function
In the case of quadratic loss function (Gaussian likelihood), it is possible to show that the predictive posterior of $\mathbf{S}_{0}=f_{\theta}(\mathbf{x}^{0})$ is a also mixture of Gaussian distributions, where the mixture is now both on the covariance matrix and on the vector of means and it is driven by an updated version of $\mathcal{Q}_{L,N}$. In the next result we assume that the matrix $\tilde{X}=[\mathbf{x}^{0},\mathbf{x}^{1},\dots,\mathbf{x}^{P}]$ is full rank.
Proposition 6
Let $\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L})$ be as in Eq. (8) and $Q^{(L)}=Q^{(L)}(Q_{1},\dots,Q_{L})$ be defined as in Proposition 1. Let $p_{\mathrm{pred}}(\mathbf{s}_{0}|\mathbf{y}_{1:P},\mathbf{x}_{0},X)$ be the posterior predictive of the FC-DLN and assume that $\det(\tilde{X}^{\top}\tilde{X})>0$. Then for every $\mathbf{s}_{0}\in\mathbb{R}^{D}$ one has
$\begin{split}&p_{\mathrm{pred}}(\mathbf{s}_{0}|\mathbf{y}_{1:P},\mathbf{x}_{0},X)\\ &\quad=\int_{(\mathcal{S}^{+}_{D})^{L}}\!\!\frac{e^{-\frac{1}{2}(\mathbf{s}_{0}-\mathbf{m}_{0})^{\top}(\Sigma_{00}-\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DP})^{-1}\Sigma_{01}^{\top})^{-1}(\mathbf{s}_{0}-\mathbf{m}_{0})}}{(2\pi)^{\frac{D}{2}}\det(\Sigma_{00}-\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DP})^{-1}\Sigma_{01}^{\top})^{\frac{1}{2}}}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\beta)\\ \end{split}$
where:
$\begin{split}&\Sigma_{00}=\frac{\mathbf{x}_{0}^{\top}\mathbf{x}_{0}}{N_{0}\lambda^{*}}\otimes Q^{(L)},\quad\Sigma_{01}=\frac{\mathbf{x}_{0}^{\top}X}{N_{0}\lambda^{*}}\otimes Q^{(L)},\quad\Sigma_{11}=\frac{X^{\top}X}{N_{0}\lambda^{*}}\otimes Q^{(L)}\\ &\mathbf{m}_{0}=\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DP})^{-1}\mathbf{y}_{1:P},\\ \end{split}$
$\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\beta):=\frac{e^{-\frac{1}{2}\Phi_{\beta}(Q_{1}\dots Q_{L},\mathbf{y}_{1:P})}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L})}{\int_{(\mathcal{S}^{+}_{D})^{L}}e^{-\frac{1}{2}\Phi_{\beta}(Q_{1}\dots Q_{L},\mathbf{y}_{1:P})}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L})}$
and $\Phi_{\beta}(Q_{1}\dots Q_{L},\mathbf{y}_{1:P}):=\mathbf{y}_{1:P}^{\top}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DP})^{-1}\mathbf{y}_{1:P}+\log(\det(\mathbb{1}_{DP}+\beta\Sigma_{11}))$.
An analogous result holds for C-DLN, where now the output is in dimension $1$. Here we need to define $\tilde{X}=[x^{\mu}_{a_{0},i}]_{\mu,a_{0},i}$ where now $\mu=0,\dots,P$ and, for every $a_{0}=1,\dots,N_{0}$, the matrix $\tilde{X}_{a_{0}}=[x^{\mu}_{a_{0},i}]_{\mu,i}$.
Proposition 7
Let $\mathcal{Q}_{L,C}(dQ_{1}\dots dQ_{L})$ be defined in Eq. (14) and $p_{\mathrm{pred}}(s_{0}|\mathbf{y}_{1:P},\mathbf{x}_{0},X)$ be the posterior predictive of the C-DLN. Assume that $\det(\sum_{a_{0}=1}^{N_{0}}\tilde{X}_{a_{0}}^{\top}\tilde{X}_{a_{0}})>0$. Then, for any $s_{0}\in\mathbb{R}$ one has
$\begin{split}&p_{\mathrm{pred}}(s_{0}|\mathbf{y}_{1:P},\mathbf{x}_{0},X)\\ &\quad=\int_{(\mathcal{S}^{+}_{N_{0}})^{L}}\frac{e^{-\frac{1}{2}(\Sigma_{00}-\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{P})^{-1}\Sigma_{01}^{\top})^{-1}(s_{0}-m_{0})^{2}}}{(2\pi)^{\frac{1}{2}}\det(\Sigma_{00}-\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{P})^{-1}\Sigma_{01}^{\top})^{\frac{1}{2}}}\mathcal{Q}_{L,C}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\beta)\\ \end{split}$
where:
$\begin{split}&\Sigma_{00}:=\sum_{r,s}\mathcal{T}(Q_{1},\dots,Q_{1})_{rs}\sum_{a_{0}}\frac{x^{0}_{a_{0},r}x^{0}_{a_{0},s}}{\lambda^{*}C_{0}},\\ &\Sigma_{01}:=\Big{[}\sum_{r,s}\mathcal{T}(Q_{1},\dots,Q_{1})_{rs}\sum_{a_{0}}\frac{x^{\mu}_{a_{0},r}x^{0}_{a_{0},s}}{\lambda^{*}C_{0}}\Big{]}_{\mu=1,\dots,P}\\ &\Sigma_{11}:=\Big{[}\sum_{r,s}\mathcal{T}(Q_{1},\dots,Q_{1})_{rs}\sum_{a_{0}}\frac{x^{\mu}_{a_{0},r}x^{\nu}_{a_{0},s}}{\lambda^{*}C_{0}}\Big{]}_{\mu,\nu=1,\dots,P},\\ &m_{0}:=\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{P})^{-1}\mathbf{y}_{1:P},\\ &\mathcal{Q}_{L,C}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\beta):=\frac{e^{-\frac{1}{2}\Phi_{\beta}(Q_{1}\dots Q_{L},\mathbf{y}_{1:P})}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L})}{\int_{(\mathcal{S}^{+}_{N_{0}})^{L}}e^{-\frac{1}{2}\Phi_{\beta}(Q_{1}\dots Q_{L},\mathbf{y}_{1:P})}\mathcal{Q}_{L,C}(dQ_{1}\dots dQ_{L})}\\ \end{split}$
and $\Phi_{\beta}(Q_{1}\dots Q_{L},\mathbf{y}_{1:P}):=\mathbf{y}_{1:P}^{\top}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{P})^{-1}\mathbf{y}_{1:P}+\log(\det(\mathbb{1}_{P}+\beta\Sigma_{11}))$.
3.3 Quantitative description of the feature learning infinite-width regime
In this section, we give a quantitative description of the feature learning infinite-width regime, using large deviation theory. For concreteness, we specialize to the case of FC-DLNs with multiple outputs, but analogous statements hold for C-DLNs.
Following the literature (Chizat et al., 2019; Mei et al., 2018; Geiger et al., 2021, 2020; Yang and Hu, 2021), we recall that the easiest way to escape the lazy-training infinite-width limit is to consider the so-called mean-field parametrization, which corresponds to the following rescaling of the loss and output functions:
$\mathcal{L}_{N}(\mathbf{y}_{1:P}|\mathbf{s}_{1:P},\beta):=\mathcal{L}(\mathbf{y}_{1:P}|\mathbf{s}_{1:P}/\sqrt{N},N\beta).$ (16)
As starting point, we briefly consider the lazy-training infinite-width limit. Here $L$ and $D$ are fixed and $N_{1}=\dots,N_{L}=N$ and $N\to+\infty$. It is easy to see that in this case $(\tilde{Q}_{1}/(\lambda_{1}N_{1}),\dots,\tilde{Q}_{L}/(\lambda_{L}N_{L}))$, converges (almost surely) to $(\mathbb{1}_{D},\dots,\mathbb{1}_{D})$. This is a consequence of the law of large numbers, indeed $Q_{\ell}=\frac{\tilde{Q}_{\ell}}{N}=\frac{1}{N}\sum_{j=1}^{N}\mathbf{Z}_{j,\ell}\mathbf{Z}_{j,\ell}^{\top}$ with $\mathbf{Z}_{j,\ell}$ independent standard Gaussian’s vectors. Since the map $(Q_{1},\dots,Q_{L})\mapsto Q^{(L)}$ in Proposition 1 is continuous, also $Q^{(L)}$ converges almost surely to $\mathbb{1}_{D}$. In this way, one recovers the well-known Gaussian lazy-training infinite-width limit.
For $L$ fixed, it is easy to derive a large deviation principle (LDP) for the random covariance matrix $Q^{(L)}$ appearing in the prior representation given in Proposition 1 or, equivalently, for $\mathcal{Q}_{L,N}$.
Roughly speaking, if $V_{N}\in\mathbb{R}^{D}$ satisfies an LDP with rate function $I$ then $P\{V_{N}\in dx\}=P_{N}(dx)\simeq e^{-NI(x)}dx$. The formal definition (see, e.g., Dembo and Zeitouni, 2010) is as follows: a sequence of random elements $(V_{N})_{N}$ taking values in a metric space $(\mathbb{V},d_{V})$ (or equivalently the sequence of their laws, say $P_{N}$) satisfies an LDP with rate function $I$ if
$\limsup_{N\to+\infty}\frac{1}{N}\log\Big{(}\mathbb{P}\{V_{N}\in C\}\Big{)}=\limsup_{N\to+\infty}\frac{1}{N}\log\Big{(}P_{N}(C)\Big{)}\leq-\inf_{v\in C}I(v)$
for any closed set $C\subset\mathbb{V}$ and
$\liminf_{N\to+\infty}\frac{1}{N}\log\Big{(}\mathbb{P}\{V_{N}\in O\}\Big{)}=\liminf_{N\to+\infty}\frac{1}{N}\log\Big{(}P_{N}(O)\Big{)}\geq-\inf_{v\in O}I(v)$
for any open set $O\subset\mathbb{V}$.
Proposition 8
The sequence of measures $\mathcal{Q}_{L,N}$, satisfies an LDP on $({\mathcal{S}^{+}_{D}})^{L}$ as $N\to+\infty$ ($L$ fixed) with rate function
$I(Q_{1},\dots,Q_{L}):=\frac{1}{2}\sum_{\ell=1}^{L}\Big{(}\operatorname{tr}(Q_{\ell})-\log(\det(Q_{\ell})\Big{)}-\frac{DL}{2}.$
Moreover, the sequence of measures $\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\beta)$ defined in Proposition 6, satisfies an LDP with the same rate function, independently from $\beta$ and $\mathbf{y}_{1:P}$.
Remark 9
As a corollary, using the well-known contraction principle (see, e.g., Theorem 4.2.1 in Dembo and Zeitouni, 2010), if $(Q_{1},\dots,Q_{L})$ has law $\mathcal{Q}_{L,N}$, then $Q^{(L)}(Q_{1},\dots,Q_{L})$ satisfies an LDP on $\mathcal{S}^{+}_{D}$ with rate function
$I^{*}(Q)=\inf\{I(Q_{1},\dots,Q_{L}):Q^{(L)}(Q_{1},\dots,Q_{L})=Q\}$
After the re-scaling in Eq. (16), and by Proposition 6, the posterior predictive distribution is a mixture of Gaussians with mixing measure
$\mathcal{Q}_{L,N}^{\circ}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\beta):=\frac{e^{-\frac{N}{2}\Phi^{\circ}_{\beta}(Q_{1},\dots,Q_{L},\mathbf{y}_{1:P})-\frac{1}{2}R(Q_{1},\dots,Q_{L})}\mathcal{Q}_{L}(Q_{1},\dots,Q_{L})}{\int_{(\mathcal{S}^{+}_{D})^{L}}e^{-\frac{N}{2}\Phi^{\circ}_{\beta}(Q_{1},\dots,Q_{L},\mathbf{y}_{1:P})-\frac{1}{2}R(Q_{1},\dots,Q_{L})}\mathcal{Q}_{L}(Q_{1},\dots,Q_{L})}$
where
$\begin{split}&\Phi_{\beta}^{\circ}(Q_{1},\dots,Q_{L},\mathbf{y}_{1:P}):=\mathbf{y}_{1:P}^{\top}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DP})^{-1}\mathbf{y}_{1:P}\\ &R(Q_{1},\dots,Q_{L})=\log(\det(\mathbb{1}_{DP}+\beta\Sigma_{11})).\\ \end{split}$ (17)
Combining well-known Varadhan’s lemma together with Proposition 8, one can prove that $\mathcal{Q}_{L,N}^{\circ}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\beta)$ also satisfies an LDP.
Proposition 10
The sequence of measures $\mathcal{Q}^{\circ}_{L,N}(dQ_{1},\dots,Q_{L}|\mathbf{y}_{1:P},\beta)$ as $N\to+\infty$ satisfies an LDP with rate function
$I^{\circ}(Q_{1},\dots,Q_{L})=\frac{1}{2}\sum_{\ell=1}^{L}\Big{(}\operatorname{tr}(Q_{\ell})-\log(\det(Q_{\ell}))\Big{)}-\frac{1}{2}\Phi_{\beta}^{\circ}(Q_{1},\dots,Q_{L},\mathbf{y}_{1:P})-\bar{I}_{0}.$
where $\bar{I}_{0}=\frac{1}{2}\inf_{Q_{1},\dots,Q_{L}}\{\sum_{\ell=1}^{L}\Big{(}\operatorname{tr}(Q_{\ell})-\log(\det(Q_{\ell}))\Big{)}-\Phi_{\beta}^{\circ}(Q_{1},\dots,Q_{L},\mathbf{y}_{1:P})\}$.
The rate function now explicitly depends on the training inputs and labels, showing that the measure over the Wishart ensembles this time concentrates around non-trivial data-dependent solutions, which quantitatively characterize feature learning in this limit (see also Discussion in Section 4 for a comparison with the proportional limit studied in Li and Sompolinsky, 2021; Pacelli et al., 2023; Hanin and Zlokapa, 2023). We observe that the result in Proposition 10 is complementary to the recent investigation in Chizat et al. (2022), where the feature learning infinite-width limit of DLNs under gradient descent dynamics was characterized.
Remark 11
Propositions 8 and  10 hold for the C-DLN with $\mathcal{Q}_{L,N}$ and $\mathcal{Q}^{\circ}_{L,N}$ replaced by $\mathcal{Q}_{L,C}$ and $\mathcal{Q}^{\circ}_{L,C}$. In this case $C\to+\infty$ and $\Phi_{\beta}^{\circ}$ has the same form given in (17) but $\Sigma_{11}$ is defined as in Proposition 7.
4 Discussion: from Kernel renormalization to rigorous Bayesian statistics
In concluding, we discuss how our derivations provide a way to interpret the recent statistical-physics literature on Bayesian deep learning theory in the proportional limit ($P,N_{\ell}\to\infty$, fixed $\alpha_{\ell}=P/N_{\ell}$) such as Li and Sompolinsky (2021, 2022); Pacelli et al. (2023); Aiudi et al. (2023).
Hanin and Zlokapa (2023) provides the first notable results that partly address this issue. However, the representation of the Bayesian model evidence (partition function) in terms of Meijer G-functions hides the possibility of a direct comparison with the previously-established physics results. On the contrary, our explicit non-asymptotic integral representation of the prior delivers a precise one-to-one mapping between the idea of kernel renormalization introduced by physicists (Li and Sompolinsky, 2021; Pacelli et al., 2023; Aiudi et al., 2023) and the Wishart ensembles that appear in our rigorous derivation.
Let us first consider the case of single output DLNs, $D=1$. Using the explicit integral representation for the prior given in Proposition 1, and assuming Gaussian likelihood, we find the following non-asymptotic result for the partition function:
$Z_{\mathbf{y}_{1:P},\beta}\propto\int_{0}^{\infty}\prod_{\ell=1}^{L}dQ_{\ell}\,e^{-\sum_{\ell=1}^{L}\left[\frac{N_{\ell}}{2}Q_{\ell}-\left(\frac{N_{\ell}}{2}-1\right)\log Q_{\ell}\right]-\frac{1}{2}\phi_{\beta}(Q_{1},\dots,Q_{L},\mathbf{y}_{1:P})}\,,$ (18)
which is exact up to an overall unimportant constant. Note that here the first two terms in the exponential arise from the probability measure over the scalar variables $Q_{\ell}$, while the third and fourth data-dependent terms are a consequence of the Gaussian integration over the output variables. This partition function can be compared with the result for non-linear networks found in the proportional limit in Pacelli et al. (2023). It turns out that if we restrict our focus to linear NNGP kernels in Pacelli et al. (2023), we find an effective action equivalent to the exponent in Eq. (18) at leading order in $N_{\ell}$. Moreover, it is reasonable to expect that the third and fourth terms in Eq. (18) scale linearly with $P$, under proper hypotheses on the training data (note that this requirement is similar to the one in Hanin and Zlokapa, 2023). This observation suggests that it is legitimate to employ large deviations techniques (or the more powerful saddle-point method) in the proportional limit, and it provides the interpretation mentioned at the beginning of this section, since one can check that the saddle-point equations derived from Eq. (18) are exactly the same as those found in Li and Sompolinsky (2021).
By a line of reasoning similar to that exposed above for $D=1$, we can also interpret the (heuristic) kernel shape renormalization pointed out in Li and Sompolinsky (2021) and Pacelli et al. (2023) for one-hidden-layer fully-connected architectures with multiple outputs, as well as the local kernel renormalization proposed to describe shallow CNNs in Aiudi et al. (2023). If we consider for instance the effective action derived in  Aiudi et al. (2023), we again recognize that the first two terms arise from the probability measure of a Wishart ensemble, whereas the last two data-dependent terms are a consequence of the integration of the output variables. Even in this case, one can check a precise equivalence with the non-asymptotic partition function that one finds in the CNN case using our result for the prior in Proposition 6 for $L=1$.
Finally, we point out that the feature learning infinite-width limit investigated in Proposition 10 shares many similarities with the proportional limit studied in the physics literature. In both cases (but for different reasons), the measure over the Wishart ensembles concentrates over data-dependent solutions of a rate function (or a saddle-point effective action). Furthermore, the three terms of the rate function in Proposition 10 also appear in the effective action found in Pacelli et al. (2023), which is known to reproduce the saddle-point equations firstly identified in Li and Sompolinsky (2021) for the linear case.
Acknowledgments and Disclosure of Funding
AP.R. is supported by ##\##NEXTGENERATIONEU (NGEU) and funded by the Ministry of University and Research (MUR), National Recovery and Resilience Plan (NRRP), project MNESYS (PE0000006) “A Multiscale integrated approach to the study of the nervous system in health and disease” (DN. 1553 11.10.2022). F.B. is partially supported by the MUR - PRIN project “Discrete random structures for Bayesian learning and prediction” no. 2022CLTYP4.
Appendix A Proofs
A.1 Preliminary facts
Matrix normal distribution. A random matrix $Z$ of dimension $n_{1}\times n_{2}$ has a centred matrix normal distribution with parameters $(\Sigma_{1},\Sigma_{2})$ (with $\Sigma_{i}$’s positive symmetric $n_{i}\times n_{i}$ matrices), if for any matrix $S$ of dimension $n_{2}\times n_{1}$
$\mathbf{E}[e^{i\operatorname{tr}(SZ)}]=\exp\left\{-\frac{1}{2}\operatorname{tr}(S\Sigma_{1}S^{\top}\Sigma_{2})\right\}.$ (19)
In symbols $Z\sim\mathcal{MN}(0,\Sigma_{1},\Sigma_{2})$.
We collect here some properties of multivariate normals, matrix normals and Wishart distributions to be used later. See Gupta and Nagar (2000) for details. In what follows $\operatorname{vec}(A)$ denotes the stacking of the columns of a matrix $A$ to form a vector.
  1. (P1)
    Linear transformation of matrix normals. Given two matrices $H$ and $K$ with compatible shape, if $Z\sim\mathcal{MN}(0,\Sigma_{1},\Sigma_{2})$ then
    $HZK\sim\mathcal{MN}(0,H\Sigma_{1}H^{\top},K^{\top}\Sigma_{2}K).$
  2. (P2)
    Equivalence with the multivariate normal. $Z\sim\mathcal{MN}(0,\Sigma_{1},\Sigma_{2})$ if and only if $\operatorname{vec}(Z)\sim\mathcal{N}(0,\Sigma_{2}\otimes\Sigma_{1})$
  3. (P3)
    Laplace functional of a Wishart distribution. Let $S$ be a symmetric positive definite matrix. Then for any symmetric matrix $C$ and any real number $\alpha$ such that all the eigenvalues of $(\mathbb{1}_{D}+\alpha SC)$ are strictly positive,
    $\begin{split}[\det(\mathbb{1}_{D}+\alpha SC)]^{-\frac{N}{2}}&=\int_{\mathcal{S}^{+}_{D}}e^{-\frac{\alpha}{2}\operatorname{tr}(CQ)}\mathcal{W}_{D}(dQ|S,N).\\ \end{split}$ (20)
    Versions of this formula can be found in literature as Ingham-Siegel integrals (Ingham, 1933; Siegel, 1935, and  Fyodorov, 2002 for an extension to Hermitian matrices); in this generality, the statement can be derived by Proposition 8.3 and its proof in Eaton (2007). In particular the formula above holds for $\alpha>0$ and $S=\mathbb{1}_{D}$ and $C$ symmetric and $\geq 0$, showing that when $\mathbf{Z}_{j}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,C)$, then
    $\mathbf{E}[e^{-\frac{\alpha}{2}\sum_{j=1}^{N}\|\mathbf{Z}_{j}\|^{2}}]=\int_{\mathcal{S}^{+}_{D}}e^{-\frac{\alpha}{2}\operatorname{tr}(CQ)}\mathcal{W}_{D}(dQ|\mathbb{1}_{D},N).$ (21)
A.2 Proofs for the FC-DLN
We set $h_{i_{\ell},\mu}^{(\ell)}=h_{i_{\ell}}^{(\ell)}(\mathbf{x}^{\mu})$ and
$H^{0}:=X=[\mathbf{x}^{1},\dots,\mathbf{x}^{P}].$
For $\ell\geq 2$ define the $N_{\ell-1}\times P$ matrix
$H^{(\ell-1)}=[h_{i,\mu}^{(\ell-1)}]_{i,\mu}.$
Notice that the weights of the network distributed as (6) can be arranged as rectangular matrices with law
$W^{(\ell-1)}\sim\mathcal{MN}(0,\mathbb{1}_{N_{\ell}},{\lambda_{\ell-1}^{-1}}\mathbb{1}_{N_{\ell-1}}).$
With this notations, it is immediate to check that for any $\ell$ one has
$H^{(\ell)}=\frac{1}{\sqrt{N_{\ell-1}}}W^{(\ell-1)}H^{(\ell-1)}\,.$ (22)
  1. (P4)
    Let $\mathbf{H}^{(\ell)}_{i_{\ell}}=(h_{i_{\ell},1}^{(\ell-1)},\dots,h_{i_{\ell},P}^{(\ell-1)})^{\top}$ the $i_{\ell}$-th row of $H^{(\ell)}$. Given $H^{(\ell-1)}$, the vectors $\mathbf{H}^{(\ell)}_{i_{\ell}}$ are independent with law $\mathcal{N}(0,(\lambda_{\ell-1}N_{\ell-1})^{-1}H^{(\ell-1)\top}H^{(\ell-1)})$. Indeed, $\mathbf{H}^{(\ell)}_{i_{\ell}}=N_{\ell-1}^{-\frac{1}{2}}W^{(\ell-1)}_{i_{\ell},\cdot}H^{(\ell-1)}$ and the $W^{(\ell-1)}_{i_{\ell},\cdot}$’s are independent and identically distributed. Recall that if $\mathbf{Z}\sim\mathcal{N}(0,C)$ and $A$ is a matrix then $A\mathbf{Z}\sim\mathcal{N}(0,ACA^{\top})$.
Proof of (10) in Remark 3. Iterating (22) one gets
$H^{(L+1)}=\frac{1}{\sqrt{N_{L}}}W^{(L)}H^{(L)}=\cdots=\frac{1}{\sqrt{N_{L}}}W^{(L)}\cdots\frac{1}{\sqrt{N_{0}}}W^{(0)}X\,.$ (23)
Proof of Proposition 1 and of (9) in Remark 3. Starting from (22) with $\ell=L+1$, conditioning on $H^{(L)}$, by (P1) and (19), it follows that for any $P\times N_{L+1}=P\times D$ matrix $\bar{S}$
$\mathbf{E}[e^{i\operatorname{tr}(H^{(L+1)}\bar{S})}]=\mathbf{E}[e^{-\frac{1}{2N_{L}\lambda_{L}}\operatorname{tr}(\bar{S}^{\top}H^{(L)\top}H^{(L)}\bar{S})}]\,.$ (24)
We divide the proof in few steps.
Step 1. For any $\ell\geq 1$ and any $P\times D$ matrix $\Theta$
$\operatorname{tr}(\Theta^{\top}H^{(\ell)\top}H^{(\ell)}\Theta)=\sum_{i_{\ell}=1}^{N_{\ell}}\|\mathbf{Y}_{i_{\ell}}^{\ell}\|^{2}$ (25)
with $\mathbf{Y}_{i_{\ell-1}}^{\ell}=H^{(\ell)}_{i_{\ell},\cdot}\Theta$. To see this, note that
$\begin{split}\operatorname{tr}(\Theta^{\top}H^{(\ell)\top}H^{(\ell)}\Theta)&=\sum_{i}\sum_{\mu,\nu}\Theta_{\mu,i}\Big{(}\sum_{i_{\ell}}h^{(\ell)}_{i_{\ell},\mu}h^{(\ell)}_{i_{\ell},\nu}\Big{)}\Theta_{\nu,i}=\sum_{i_{\ell}}\sum_{i}\sum_{\mu,\nu}\Theta_{\mu,i}\Theta_{\nu,i}h^{(\ell)}_{i_{\ell},\mu}h^{(\ell)}_{i_{\ell},\nu}\\ &=\sum_{i_{\ell}}\sum_{i}\Big{(}\sum_{\mu}h^{(\ell)}_{i_{\ell},\mu}\Theta_{\mu,i}\Big{)}^{2}=\sum_{i_{\ell}}\|\mathbf{Y}_{i_{\ell}}^{\ell}\|^{2}.\\ \end{split}$
Step 2. For any $\ell\geq 1$ and any $P\times D$ matrix $\Theta$ we claim that
$\mathbf{E}[e^{-\frac{1}{2\lambda_{\ell}N_{\ell}}\operatorname{tr}(\Theta^{\top}H^{(\ell)\top}H^{(\ell)}\Theta)}]=\mathbf{E}[e^{-\frac{1}{2\lambda_{\ell}}\operatorname{tr}(K_{\ell-1}Q_{\ell})}]$ (26)
where $Q_{\ell}$ is a $D\times D$ random matrix with Wishart distribution of $N_{\ell}$ degree of freedom and scale matrix $\mathbb{1}_{D}/N_{\ell}$ and
$K_{\ell-1}:=\frac{1}{\lambda_{\ell-1}N_{\ell-1}}\Theta^{\top}H^{(\ell-1)\top}H^{(\ell-1)}\Theta.$
Writing $Q_{\ell}=U_{\ell}^{\top}U_{\ell}$, with $U_{\ell}$ a $D\times D$ matrix, one has
$\mathbf{E}[e^{-\frac{1}{2\lambda_{\ell}N_{\ell}}\operatorname{tr}(\Theta^{\top}H^{(\ell)\top}H^{(\ell)}\Theta)}]=\mathbf{E}[e^{-\frac{1}{2\lambda_{\ell}N_{\ell-1}\lambda_{\ell-1}}\operatorname{tr}(U_{\ell}\Theta^{\top}H^{(\ell-1)\top}H^{(\ell-1)}\Theta U_{\ell-1}^{\top})}].$ (27)
To prove (26), note that by (P4) the $\mathbf{Y}_{i_{\ell}}^{\ell}$’s are independent and identically distributed conditionally on $H^{\ell-1}$ , with
$\mathbf{Y}_{i_{\ell}}^{\ell}|H^{(\ell-1)}\sim\mathcal{N}(0,K_{\ell-1}).$
Hence using, (25) and (21)
$\mathbf{E}[e^{-\frac{1}{2\lambda_{\ell}N_{\ell}}\operatorname{tr}(\Theta^{\top}H^{(\ell)\top}H^{(\ell)}\Theta)}]=\mathbf{E}[e^{-\frac{1}{2\lambda_{\ell}N_{\ell}}\sum_{i_{\ell}=1}^{N_{\ell}}\|\mathbf{Y}_{i_{\ell}}^{\ell}\|^{2}}]=\mathbf{E}[e^{-\frac{1}{2\lambda_{\ell}}\operatorname{tr}(K_{\ell-1}Q_{\ell})}].$ (28)
Step 3. Starting from $\ell=L$ and $\Theta=\bar{S}$, iterating (27) one obtains
$\mathbf{E}[e^{-\frac{1}{2\lambda_{L}N_{L}}\operatorname{tr}(\bar{S}^{\top}H^{(L)\top}H^{(L)}\bar{S})}]=\mathbf{E}[e^{-\frac{1}{2\lambda^{*}N_{0}}\operatorname{tr}(U_{1:L}\bar{S}^{\top}H^{(0)\top}H^{(0)}\bar{S}U_{1:L}^{\top})}]$
where $\lambda^{*}:=\lambda_{0}\dots\lambda_{L}$ and $U_{1:L}=U_{1}\dots U_{L}$ for $Q_{\ell}=U_{\ell}^{\top}U_{\ell}$ independent $D\times D$ random matrix with Wishart distribution with $N_{\ell}$ degree of freedom and $\mathbb{1}_{D}/N_{\ell}$ as scale matrix. By (24) we conclude that
$\begin{split}\mathbf{E}[e^{i\operatorname{tr}(H^{(L+1)}\bar{S})}]&=\mathbf{E}[e^{-\frac{1}{2\lambda_{L}N_{L}}\operatorname{tr}(\bar{S}^{\top}H^{(L)\top}H^{(L)}\bar{S})}]\\ &=\mathbf{E}[e^{-\frac{1}{2\lambda^{*}N_{0}}\operatorname{tr}(U_{1:L}\bar{S}^{\top}H^{(0)\top}H^{(0)}\bar{S}U_{1:L}^{\top})}]\\ &=\mathbf{E}[e^{-\frac{1}{2\lambda^{*}N_{0}}\operatorname{tr}(\bar{S}U_{1:L}^{\top}U_{1:L}\bar{S}^{\top}H^{(0)\top}H^{(0)})}]\,.\end{split}$
Recalling that $H^{(0)\top}H^{(0)}:=X^{\top}X$, using once again (19), this shows that, conditionally on $U_{1:L}^{\top}U_{1:L}$, $H^{(L+1)}\sim\mathcal{MN}(0,U_{1:L}^{\top}U_{1:L},(\lambda^{*}N_{0})^{-1}X^{\top}X)$. Note that this is (9) of Remark 3.
Since $\mathbf{S}_{1:P}=\operatorname{vec}[H^{(L+1)}]$, by (P2), one gets
$\mathbf{S}_{1:P}|U_{1:L}^{\top}U_{1:L}\sim\mathcal{N}(0,(\lambda^{*}N_{0})^{-1}X^{\top}X\otimes U_{1:L}^{\top}U_{1:L}).$
and the thesis follows.
$\blacklozenge$
We now prove a slightly more general statement of the one given in Proposition 6. Let $\mathbf{s}^{\top}=((\mathbf{s}^{0})^{\top},(\mathbf{s}^{1})^{\top},\dots,(\mathbf{s}^{P})^{\top})$ and $\tilde{X}=[\mathbf{x}^{0},\mathbf{x}^{1},\dots,\mathbf{x}^{P}]$. In what follows $\mathbb{1}_{M}$ is the identity matrix of dimension $M\times M$ and $0_{M\times N}$ is the zero matrix of dimension $M\times N$.
Proposition 12
Let $\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L})$ be as in Eq. (8) and $Q^{(L)}=Q^{(L)}(Q_{1},\dots,Q_{L})$ be defined as in Proposition 1. Let $p_{\mathrm{post}}(\mathbf{s}|\mathbf{y}_{1:P},\mathbf{x}_{0},X)$ be the posterior distribution of $(\mathbf{S}_{0},\dots,\mathbf{S}_{P})$ given $\mathbf{y}_{1:P}$ in a FC-DLN. If $\det(\tilde{X}^{\top}\tilde{X})>0$, then
$p_{\mathrm{post}}(\mathbf{s}|\mathbf{y}_{1:P},\tilde{X})\\ =\int_{(\mathcal{S}^{+}_{D})^{L}}\frac{e^{-\frac{1}{2}(\mathbf{s}-\mathbf{m})^{\top}(\beta\Pi_{0}+\Sigma^{-1})(\mathbf{s}-\mathbf{m})}}{(2\pi)^{\frac{D(P+1)}{2}}\det((\beta\Pi_{0}+\Sigma^{-1})^{-1})^{\frac{1}{2}}}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\tilde{X},\beta)$
where
$\Sigma=\begin{pmatrix}\Sigma_{00}&\Sigma_{01}\\ \Sigma_{01}^{\top}&\Sigma_{11}\\ \end{pmatrix}:=\frac{\tilde{X}^{\top}\tilde{X}}{N_{0}\lambda^{*}}\otimes Q^{(L)},\qquad\Pi_{0}=\begin{pmatrix}0_{D\times D}&0_{P\times DP}\\ 0_{DP\times P}&\mathbb{1}_{DP}\\ \end{pmatrix}$
$\mathbf{m}:=\begin{pmatrix}\mathbf{m}_{0}\\ \mathbf{m}_{1}\\ \end{pmatrix}=\begin{pmatrix}\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DP})^{-1}\\ \Sigma_{11}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DP})^{-1}\\ \end{pmatrix}\mathbf{y}_{1:P},$
$\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\tilde{X},\beta):=\frac{e^{-\frac{1}{2}\Phi_{\beta}(Q_{1},\dots,Q_{L},\mathbf{y}_{1:P})}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L})}{\int_{(\mathcal{S}^{+}_{D})^{L}}e^{-\frac{1}{2}\Phi_{\beta}(Q_{1},\dots,Q_{L},\mathbf{y}_{1:P})}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L})}$
and $\Phi_{\beta}(Q_{1},\dots,Q_{L},\mathbf{y}_{1:P}):=\mathbf{y}_{1:P}^{\top}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DK})^{-1}\mathbf{y}_{1:P}+\log(\det(\mathbb{1}_{DP}+\beta\Sigma_{11}))$.
Proof.    Proposition 1 gives
$\mathcal{L}(\mathbf{y}_{1:P}|\mathbf{s}_{1:P})p_{\mathrm{prior}}(\mathbf{s}_{0},\mathbf{s}_{1:P}|\mathbf{x}^{0},X)=\int\frac{e^{-\frac{\beta}{2}(\mathbf{s}_{1:P}-\mathbf{y}_{1:P})^{\top}(\mathbf{s}_{1:P}-\mathbf{y}_{1:P})-\frac{1}{2}\mathbf{s}\Sigma^{-1}\mathbf{s}}}{\det(\Sigma)^{\frac{1}{2}}(2\pi)^{\frac{D(P+1)}{2}}}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L})$
with $\Sigma:=(N_{0}\lambda^{*})^{-1}\tilde{X}^{\top}\tilde{X}\otimes Q^{(L)}$. Since $\tilde{X}^{\top}\tilde{X}$ is assumed to be strictly positive definite and hence invertible, using that $(A\otimes B)^{-1}=A^{-1}\otimes B^{-1}$ (see, e.g., Proposition 1.28 Eaton, 2007) it follows that also $\Sigma$ is invertible. Simple computations show that
$\begin{split}{\beta}(\mathbf{s}_{1:P}-\mathbf{y}_{1:P})^{\top}(\mathbf{s}_{1:P}-\mathbf{y}_{1:P})+\mathbf{s}\Sigma^{-1}\mathbf{s}&=(\mathbf{s}-\tilde{\mathbf{y}})^{\top}{\Pi_{0}}{\beta}(\mathbf{s}-\tilde{\mathbf{y}})+\mathbf{s}\Sigma^{-1}\mathbf{s}\\ &=(\mathbf{s}-\mathbf{m})^{\top}({\Pi_{0}}{\beta}+\Sigma^{-1})(\mathbf{s}-\mathbf{m})+\varphi(\Sigma,\mathbf{y}_{1:P})\\ \end{split}$
where
$\begin{split}&\tilde{\mathbf{y}}:=(0_{D}^{\top},\mathbf{y}_{1:P}^{\top})^{\top},\quad\mathbf{m}=({\Pi_{0}}{\beta}+\Sigma^{-1})^{-1}{\Pi_{0}}{\beta}\tilde{\mathbf{y}}={\beta}({\Pi_{0}}{\beta}+\Sigma^{-1})^{-1}\tilde{\mathbf{y}}\,,\\ &\varphi(\Sigma,\mathbf{y}_{1:P})=\beta{\mathbf{y}_{1:P}^{\top}\mathbf{y}_{1:P}}-(\mathbf{m})^{\top}({\Pi_{0}}{\beta}+\Sigma^{-1})\mathbf{m}\,.\\ \\ \end{split}$
Writing $\mathbf{m}:=[\mathbf{m}_{0},\mathbf{m}_{1}]$ one can check that $(\Sigma_{11}^{-1}+\beta\mathbb{1}_{DP})^{-1}=\Sigma_{11}(\mathbb{1}_{DP}+\beta\Sigma_{11})^{-1}$ and
$\begin{split}&\mathbf{m}_{1}=\beta(\Sigma_{11}^{-1}+\beta\mathbb{1}_{DP})^{-1}\mathbf{y}_{1:P}=\beta\Sigma_{11}(\mathbb{1}_{DP}+\beta\Sigma_{11})^{-1},\\ &\mathbf{m}_{0}=\beta\Sigma_{01}\Sigma_{11}^{-1}(\Sigma_{11}^{-1}+\beta\mathbb{1}_{DP})^{-1}\mathbf{y}_{1:P}=\beta\Sigma_{01}(\mathbb{1}_{DP}+\beta\Sigma_{11})^{-1}\mathbf{y}_{1:P},\\ &\varphi(\Sigma,\mathbf{y}_{1:P})=\beta{\mathbf{y}_{1:P}^{\top}\mathbf{y}_{1:P}}-\beta^{2}\mathbf{y}_{1:P}^{\top}(\Sigma_{11}^{-1}+\beta\mathbb{1}_{DP})^{-1}\mathbf{y}_{1:P}=\mathbf{y}_{1:P}^{\top}(\Sigma_{11}^{-1}+\mathbb{1}_{DP}/\beta)^{-1}\mathbf{y}_{1:P}.\\ \end{split}$
In conclusion
$\frac{e^{-\frac{\beta}{2}(\mathbf{s}_{1:P}-\mathbf{y}_{1:P})^{\top}(\mathbf{s}_{1:P}-\mathbf{y}_{1:P})-\frac{1}{2}\mathbf{s}\Sigma^{-1}\mathbf{s}}}{|\Sigma|^{\frac{1}{2}}}=\frac{e^{-\frac{1}{2}\varphi(\Sigma,\mathbf{y}_{1:P})}}{\det(\Sigma)^{\frac{1}{2}}\det(\beta{\Pi_{0}}+\Sigma^{-1})^{\frac{1}{2}}}\frac{e^{-\frac{1}{2}(\mathbf{s}-\mathbf{m})^{\top}(\beta\Pi_{0}+\Sigma^{-1})(\mathbf{s}-\mathbf{m})}}{(\det(\beta\Pi_{0}+\Sigma^{-1})^{-1})^{\frac{1}{2}}}.$
where $\det(\Sigma)^{\frac{1}{2}}\det(\beta{\Pi_{0}}+\Sigma^{-1})^{\frac{1}{2}}=\det(\mathbb{1}_{DP}+\beta\Sigma_{11})^{\frac{1}{2}}.$blacklozenge
Proof of Proposition 6. Applying the previous proposition, one gets that
[…]
is again a mixture of gaussians with mean $\mathbf{m}_{0}=\Sigma_{01}(\Sigma_{11}+\beta^{-1}I)^{-1}\mathbf{y}_{1:P}$ and (conditional) covariance
$\Big{[}\Big{(}\beta\Pi_{0}+\Sigma^{-1}\Big{)}^{-1}\Big{]}_{1:D,1:D}.$
To conclude it remains to compute the previous matrix. We have already noted that, by $\tilde{X}^{\top}\tilde{X}>0$, it follows that also $\Sigma$ is invertible. Being $X^{\top}X$ and $\mathbf{x}_{0}^{\top}\mathbf{x}_{0}$ strictly positive, then the Schur complement of $\Sigma_{00}$ in $\Sigma$, $\Sigma^{*}_{00}:=\Sigma_{00}-\Sigma_{01}\Sigma_{11}^{-1}\Sigma_{01}^{\top}$, is well-defined and strictly positive (Proposition 1.34 Eaton, 2007). In particular, using the Banachiewicz inversion formula, one can write the inverse of $\Sigma$ as
$\Sigma^{-1}=\begin{pmatrix}(\Sigma^{*}_{00})^{-1}&-(\Sigma^{*}_{00})^{-1}\Sigma_{01}\Sigma_{11}^{-1}\\ -((\Sigma^{*}_{00})^{-1}\Sigma_{01}\Sigma_{11}^{-1})^{\top}&\Sigma_{11}^{-1}+\Sigma_{11}^{-1}\Sigma_{01}^{\top}(\Sigma^{*}_{00})^{-1}\Sigma_{01}\Sigma_{11}^{-1}\\ \end{pmatrix}.$
See, e.g., Theorem 1.2 in Zhang (2005). Using again the Banachiewicz inversion formula on $\beta\Pi_{0}+\Sigma^{-1}$ (which is again strictly positive), after some tedious computations, one gets that
$\Big{[}\Big{(}\beta\Pi_{0}+\Sigma^{-1}\Big{)}^{-1}\Big{]}_{1:D,1:D}=\Sigma_{00}-\Sigma_{01}\Big{(}\frac{1}{\beta}\mathbb{1}_{DP}+\Sigma_{11}\Big{)}^{-1}\Sigma_{01}^{\top}.$
Observing that
$\Sigma_{00}=\frac{\mathbf{x}_{0}^{\top}\mathbf{x}_{0}}{N_{0}\lambda^{*}}\otimes Q^{(L)},\quad\Sigma_{01}=\frac{\mathbf{x}_{0}^{\top}X}{N_{0}\lambda^{*}}\otimes Q^{(L)},\quad\Sigma_{11}=\frac{X^{\top}X}{N_{0}\lambda^{*}}\otimes Q^{(L)},$
the proof of Proposition 6 is concluded. $\blacklozenge$
Proof of Proposition 8. We give the proof in the case $L=1$, the general case follows the same lines, or using Exercise 4.2.7 in Dembo and Zeitouni (2010) and the independence of the components. The result is a consequence of a very general version of the Cramér’s theorem. See Theorem 6.1.3 and Corollary 6.1.6 in Dembo and Zeitouni (2010). Recall that if $Q_{1}\sim\mathcal{W}_{D}(\cdot|\mathbb{1}_{D},N)$ then $Q_{\ell}=\frac{1}{N}\sum_{j=1}^{N}\mathbf{Z}_{j}\mathbf{Z}_{j}^{\top}$ where $\mathbf{Z}_{j}\mathbf{Z}_{j}^{\top}\sim\mathcal{W}_{D}(\cdot|\mathbb{1}_{D},1)$ are independent and identically distributed taking values in $\mathcal{S}^{+}_{D}$. Now $\mathcal{S}^{+}_{D}$ is a convex open cone in the (topological) vector space $\mathcal{S}_{D}$ of the $D\times D$ symmetric matrices, endowed with the scalar product $(A,B)=\operatorname{tr}(AB)$. The space $\mathcal{S}_{D}$ is locally convex, Hausdorff and the closure of $\mathcal{S}^{+}_{D}$ in $\mathcal{S}_{D}$ is convex and separable and complete (with the topology induced by $\mathcal{S}_{D})$. Clearly $P\{\mathbf{Z}_{j}\mathbf{Z}_{j}^{\top}\in\overline{\mathcal{S}^{+}_{D}}\}=1$ and any closed convex hull of a compact in $\overline{\mathcal{S}^{+}_{D}}$ is compact. Hence all the hypotheses of the General Cramér’s theorem given in Chapter 6 of Dembo and Zeitouni (2010) are satisfied and hence $Q_{\ell}$ satisfies a (weak) LDP (when $N\to+\infty$) with rate function given by $I(Q)=\Lambda^{*}(Q):=\sup_{A\in\mathcal{S}_{D}}\{\operatorname{tr}(QA)-\Lambda(A)\}$ where
$\Lambda(A)=\log\left(\int_{\mathcal{S}^{+}_{D}}e^{\operatorname{tr}(AQ)}P\{\mathbf{Z}_{1}\mathbf{Z}_{1}^{\top}\in dQ\}\right).$
In point of fact, see Corollary 6.1.6 in Dembo and Zeitouni (2010), it holds also a strong LDP provided that 00 is in the interior of the domain of $\Lambda$, as we shall show. The well-known expression for the Laplace transform of a Wishart distribution with one degree of freedom (see (20)) gives $\Lambda(A)=-\frac{1}{2}\log(\det(\mathbb{1}_{D}-2A))$ wherever the eigenvalues of $\mathbb{1}_{D}-2A$ are positive and +$\infty$+ ∞ otherwise. In computing $\Lambda^{*}(Q)$ we can restrict $A$ to be such that the eigenvalues of $\mathbb{1}_{D}-2A$ are positive, since otherwise $-\Lambda(A)=-\infty$. Now we check that
$\Lambda^{*}(Q)=\sup_{A}\{\operatorname{tr}(QA)+\frac{1}{2}\log\det(\mathbb{1}_{D}-2A)\}=\frac{1}{2}(\operatorname{tr}(Q)+\log(\det(Q^{-1}))-D).$
Making the change of variables $B=\mathbb{1}_{D}-2A$, we compute $\sup_{B}H(B,Q)$ for
$H(B,Q)=\operatorname{tr}(Q(\mathbb{1}_{D}-B)/2)+\frac{1}{2}\log(\det(B))$
where now the $\sup$ is over symmetric strictly positive matrices $B$. The gradient of $B\mapsto H(B,Q)$ is easily computed with the Jacobi’s formula, $\partial_{B_{ij}}\Lambda(A)=-Q_{ij}+(B^{-1})_{ij}$ for $i<j$ and $\partial_{B_{ii}}\Lambda(A)=-Q_{ii}/2+(B^{-1})_{ii}/2$ for $i=j$. Solving $\nabla H(B_{*},Q)=0$ we get $0=-Q+B_{*}^{-1}$ and hence $B_{*}=Q^{-1}$. Since $B\mapsto-H(B,Q)$ is convex and $-H(B,Q)\to+\infty$ when $B$ converges to the boundary of the convex cone $\mathcal{S}^{+}_{D}$ (since in this case $\log\det(B)\to-\infty$), the unique critical point $B_{*}$ is also the unique minimum. Hence $\Lambda^{*}(Q)=H(B_{*},Q)=\frac{1}{2}(\operatorname{tr}(Q)+\log(|Q^{-1}|)-D)$, as desired. The LDP for $\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\beta)$ follows now using the next Lemma 13. Details are omitted since very similar to those given in the next proof of Proposition 10. $\blacklozenge$
In order to prove Proposition 10, we need the following variant of the Varadhan’s lemma.
Lemma 13
Let ${P}_{N}$ be a sequence of probability measure on a convex closed subset $\mathcal{S}\subset\mathbb{R}^{D}$, satisfying an LDP with rate function $I$. Assume that
  • $\Phi_{0}$ is a bouned continuous function from $\mathcal{S}\to\mathbb{R}$;
  • $R:\mathcal{S}\to[0,+\infty)$ such that $\sup_{O\cap B_{R}}|R(s)|<+\infty$ for any $R$, where $B_{R}=\{s\in\mathbb{R}^{D}:\|s\|<R\}$;
  • the rate function $I$ is such that $\inf_{s\in\mathcal{S}\cap B_{R}^{c}}I(s)\to+\infty$ as $R\to+\infty$ and $\inf_{s}I(s)<+\infty$.
Then
$P_{N}^{\circ}(ds)=\frac{e^{-(N\Phi_{0}(s)+R(s))}P_{N}(ds)}{\int_{\mathcal{S}}e^{-(N\Phi_{0}(s)+R(s))}P_{N}(ds)}$
satisfies an LDP with rate function $I(s)-\Phi_{0}(s)-I_{0}$ where where $I_{0}=\inf_{s}[I(s)-\Phi_{0}(s)]$.
Proof.    Varhadan Lemma gives that $\bar{P}_{N}(ds)=P_{N}(ds)/(\int_{\mathcal{S}}e^{-N\Phi_{0}(s)}P_{N}(ds))$ satifies an LDP with rate $I_{\Phi_{0}}(s)=I(s)-\Phi_{0}(s)-I_{0}$. Then,
$\frac{1}{N}\log\Big{(}\int_{\mathcal{S}}e^{-(N\Phi_{0}(s)+R(s))}\bar{P}_{N}(ds)\Big{)}\leq\frac{1}{N}\log\Big{(}\int_{\mathcal{S}}e^{-N\Phi_{0}(s)}\bar{P}_{N}(ds)\Big{)}=0$
and
$\begin{split}\frac{1}{N}&\log\Big{(}\int_{\mathcal{S}}e^{-(N\Phi_{0}(s)+R(s))}\bar{P}_{N}(ds)\Big{)}\geq\frac{1}{N}\log\Big{(}\int_{B_{R}\cap\mathcal{S}}e^{-N\Phi_{0}(s)}\bar{P}_{N}(ds)e^{-\sup_{B_{R}\cap\mathcal{S}}|R(s)|}\Big{)}\\ &\geq-\inf_{s\in B_{R}\cap\mathcal{S}}[I(s)-\Phi_{0}(s)]-I_{0}.\\ \end{split}$
Since for $R\to+\infty$ one has $\inf_{s\in B_{R}\cap\mathcal{S}}[I(s)-\Phi_{0}(s)]-I_{0}\to 0$, one obtains that
$\frac{1}{N}\log\Big{(}\int_{\mathcal{S}}e^{-(N\Phi_{0}(s)+R(s))}\bar{P}_{N}(ds)\Big{)}\to 0$
which means that
$\frac{1}{N}\log\Big{(}\frac{\int_{\mathcal{S}}e^{-(N\Phi_{0}(s)+R(s))}P_{N}(ds)}{\int_{\mathcal{S}}e^{-N\Phi_{0}(s)}P_{N}(ds)}\Big{)}\to 0.$
If now $C$ is a closed set, then
$\begin{split}&\limsup_{N\to+\infty}\frac{1}{N}\log\Big{(}\int_{C}e^{-(N\Phi_{0}(s)+R(s))]}\bar{P}_{N}(ds)\Big{)}\leq\limsup_{N\to+\infty}\frac{1}{N}\log\Big{(}\int_{C}e^{-N\Phi_{0}(s)}\bar{P}_{N}(ds)\Big{)}\\ &\leq-\inf_{s\in C}[I(s)-\Phi_{0}(s)-I_{0}]\\ \end{split}$
On the other hand if $O$ is open, then
$\begin{split}&\liminf_{N\to+\infty}\frac{1}{N}\log\Big{(}\int_{O}e^{-(N\Phi_{0}(s)+R(s))]}\bar{P}_{N}(ds)\log\Big{)}\\ &\geq\liminf_{N\to+\infty}\frac{1}{N}\Big{[}\log\Big{(}\int_{O\cap B_{R}}e^{-N\Phi_{0}(s)}\bar{P}_{N}(ds)\log\Big{)}+\log\Big{(}e^{-\sup_{O\cap B_{R}}|R(s)|}\Big{)}\Big{]}\\ &\geq-\inf_{s\in O\cap B_{R}}[I(s)-\Phi_{0}(s)-I_{0}].\\ \end{split}$
Since $\inf_{s\in\mathcal{S}\cap B_{R}^{c}}I(s)\to+\infty$ as $R\to+\infty$, $\inf_{s}I(s)<+\infty$ and $\Phi_{0}$ is bounded $\inf_{s\in\mathcal{S}\cap B_{R}^{c}}[I(s)-\Phi_{0}(s)]\to+\infty$ as $R\to+\infty$, which means that for $R$ big enough $\inf_{s\in O\cap B_{R}}[I(s)-\Phi_{0}(s)-I_{0}]=\inf_{s\in O}[I(s)-\Phi_{0}(s)-I_{0}]$. Combining these facts the thesis follows. $\blacklozenge$
Proof. Proposition 10    The proof follows by Lemma  13. To check the assumptions, note that
$Q_{1},\dots,Q_{L}\mapsto\Sigma(Q_{1},\dots,Q_{L})=\begin{pmatrix}\Sigma_{00}&\Sigma_{01}\\ \Sigma_{01}^{\top}&\Sigma_{11}\\ \end{pmatrix}:=\frac{\tilde{X}^{\top}\tilde{X}}{N_{0}\lambda^{*}}\otimes Q^{(L)}$
is a continuous function. Moreover, $\Sigma_{11}\mapsto\mathbf{y}_{1:P}^{\top}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DP})^{-1}\mathbf{y}_{1:P}$ is continuous and bounded (on $\mathcal{S}^{+}$), and hence the same holds for
$Q_{1},\dots,Q_{L}\mapsto\mathbf{y}_{1:P}^{\top}(\Sigma_{11}(Q_{1},\dots,Q_{L})+\beta^{-1}\mathbb{1}_{DP})^{-1}\mathbf{y}_{1:P}.$
Let $\lambda_{i}$’s be the eigenvalues of $\beta\Sigma_{11}$, then $\mathbb{1}_{DP}+\beta\Sigma_{11}$ has eigenvalues $1+\lambda_{i}$ and
$\log(\det(\mathbb{1}_{DP}+\beta\Sigma_{11}))=\sum_{i=1}^{DP}\log(1+\lambda_{i}).$
hence $\Sigma_{11}\mapsto\log(\det(\mathbb{1}_{DP}+\beta\Sigma_{11}))$ is bounded on bounded set of $\Sigma_{11}$. Using the continuity of $(Q_{1},\dots,Q_{L})\mapsto\Sigma_{11}(Q_{1},\dots,Q_{L})$ it follows that also $(Q_{1},\dots,Q_{L})\mapsto\log(\det(\mathbb{1}_{DP}+\beta\Sigma_{11}))$ is bounded when $\sum_{\ell=1}^{L}\|Q_{\ell}\|_{2}\leq R$ for any $R$. It remains to check that
$\lim_{R\to+\infty}\inf_{(Q_{1},\dots,Q_{L}):\sum\|Q_{\ell}\|_{2}^{2}>R^{2}}\sum_{\ell=1}^{L}\Big{(}\operatorname{tr}(Q_{\ell})-\log(\det(Q_{\ell}))\Big{)}=+\infty.$
Letting $\lambda_{i\ell}$ the eigenvalues of $Q_{\ell}$, the claim follows since $\operatorname{tr}(Q_{\ell})-\log(\det(Q_{\ell}))=\sum_{i=1}^{P}\Big{(}\lambda_{i\ell}+\log(\lambda_{i\ell})\Big{)}$. $\blacklozenge$
A.3 Proofs for the C-DLN
Lemma 14
Let $\mathcal{K}_{\mathrm{C}}(Q_{1},\dots,Q_{L})$ be defined as in (15). Then, if $Q_{1},\dots,Q_{L}$ are strictly positive matrices, then $\mathcal{K}_{\mathrm{C}}(Q_{1},\dots,Q_{L})$ is positive. If in addition $\sum_{a_{0}}X_{a_{0}}^{\top}X_{a_{0}}$ is strictly positive, then $\mathcal{K}_{\mathrm{C}}(Q_{1},\dots,Q_{L})$ is strictly positive.
Proof. First of all note that if $Q_{1},\dots,Q_{L}$ are strictly positive then $\mathcal{T}(Q_{1},\dots,Q_{L})$ is strictly positive. For $\bar{\mathbf{s}}_{1:P}\not=0$ write
$\begin{split}\bar{\mathbf{s}}_{1:P}^{\top}\mathcal{K}_{\mathrm{C}}(Q_{1},\dots,Q_{L})\bar{\mathbf{s}}_{1:P}&=\sum_{\mu,\nu}\sum_{a_{0}}\sum_{r,s}\mathcal{T}(Q_{1},\dots,Q_{L})_{rs}\frac{x^{\mu}_{a_{0},r}x^{\nu}_{a_{0},s}}{\lambda^{*}C_{0}}\bar{s}_{\mu}\bar{s}_{\nu}\\ &=\sum_{a_{0}}\sum_{r,s}\mathcal{T}(Q_{1},\dots,Q_{L})_{rs}A^{*}_{r,a_{0}}A^{*}_{s,a_{0}}\end{split}$
with $A^{*}_{r,a_{0}}=\sum_{\mu}\bar{s}_{\mu}\frac{x^{\mu}_{a_{0},r}}{\sqrt{\lambda^{*}C_{0}}}$. Using $\mathcal{T}(Q_{1},\dots,Q_{L})>0$, one gets
$\sum_{r,s}\mathcal{T}(Q_{1},\dots,Q_{L})_{rs}A^{*}_{r,a_{0}}A^{*}_{s,a_{0}}\geq 0$
for every $a_{0}$, so that $\bar{\mathbf{s}}_{1:P}^{\top}\mathcal{K}_{\mathrm{C}}(Q_{1},\dots,Q_{L})\bar{\mathbf{s}}_{1:P}\geq 0$. If now $0=\bar{\mathbf{s}}_{1:P}^{\top}\mathcal{K}_{\mathrm{C}}(Q_{1},\dots,Q_{L})\bar{\mathbf{s}}_{1:P}$, then $A^{*}_{r,a_{0}}=0$ for every $r$ and $a_{0}$. Hence,
$0=\sum_{r}(A^{*}_{r,a_{0}})^{2}=\sum_{\mu,\nu}\bar{s}_{\mu}\bar{s}_{\nu}\sum_{r}\frac{x^{\mu}_{a_{0},r}x^{\nu}_{a_{0},r}}{{\lambda^{*}C_{0}}}$
If $\sum_{a_{0}}X_{a_{0}}^{\top}X_{a_{0}}>0$, this is possible only if $\bar{\mathbf{s}}_{1:P}=0$blacklozenge
Let $K=[K_{i,j,\mu,\nu}]_{i,j,\mu,\nu}$ a four index tensor with $i,j=1,\dots,N_{0}$ and $\mu,\nu=1,\dots,P$ corresponding to a covariance operator. One can identifies $K$ with a symmetric and positive definite $(N_{0}\times P)\times(N_{0}\times P)$ matrix $K=[K_{i,j,\mu,\nu}]_{(i,\mu),(j,\nu)}$, where the multi-indices $(i,\mu)$’s are properly ordered from $1$ to $N_{0}\times P$. With a slight abuse of language, we will use in the following $K^{(\ell-1)}$ to denote either the four indices tensor or its matrix representation, the case being clear from the context.
Lemma 15
Let $\mathbf{s}$ be a vector in $\mathbb{R}^{P}$, then
$\det[\mathbb{1}_{N_{0}}\otimes\mathbb{1}_{P}+[\mathbb{1}_{N_{0}}\otimes(\mathbf{s}\mathbf{s}^{\top})]K]$ $=\det\Big{(}\mathbb{1}_{N_{0}}+\sum_{\mu,\nu}{s}_{\mu}{s}_{\nu}K_{\cdot,\mu\nu}\Big{)}\,,$ (29)
where $K_{\cdot,\mu\nu}=[K_{i,j,\mu,\nu}]_{i,j=1}^{N_{0}}$.
Proof. The eigenvalues of $[\mathbb{1}_{N_{0}}\otimes(\mathbf{s}\mathbf{s}^{\top})]K$ are given by
$[\mathbb{1}_{N_{0}}\otimes(\mathbf{s}\mathbf{s}^{\top})]Kv=\lambda v$
which reads, in components, as
$\sum_{\mu,\nu,j}s_{\mu}K_{(i,\mu)(j,\nu)}v_{(j,\nu)}s_{\rho}=\lambda v_{(i,\rho)}\,.$
This shows that if $\lambda\not=0$ then $v_{(i,\mu)}=c_{i}s_{\mu}$ for some $\mathbf{c}\in\mathbb{R}^{N_{0}}$, such that
$\sum_{j}\Biggl{(}\sum_{\mu,\nu}s_{\mu}K_{\cdot,\mu\nu}s_{\nu}\Biggr{)}_{ij}c_{j}=\lambda c_{i}\,,$
that is $\lambda\in\operatorname{Sp}(\sum_{\mu,\nu}s_{\mu}K_{\cdot,\mu\nu}s_{\nu})$, where $\operatorname{Sp}$ denotes the spectrum. The cardinality of this spectrum is $N_{0}$, all the other $N_{0}(P-1)$ eigenvalues must be zero. Hence,
$\operatorname{Sp}(\mathbb{1}_{N_{0}}\otimes\mathbb{1}_{P}+[\mathbb{1}_{N_{0}}\otimes(\mathbf{s}\mathbf{s}^{\top})]K)=\operatorname{Sp}\Bigl{(}\mathbb{1}_{N_{0}}+\sum_{\mu,\nu}s_{\mu}K_{\cdot,\mu\nu}s_{\nu}\Bigr{)}\cup\{1\}^{N_{0}(P-1)}\,,$
and the thesis follows. $\blacklozenge$
Proof of Proposition 5.
From Eq. (5), we have
$\mathbf{E}[e^{i\bar{\mathbf{s}}_{1:P}^{\top}\mathbf{S}_{1:P}}]=\mathbf{E}\left[\exp\left(-\frac{1}{2}\sum_{\mu,\nu}\bar{s}^{\mu}\bar{s}^{\nu}\sum_{a_{L}=1}^{C_{L}}\sum_{i=1}^{N_{0}}\frac{1}{\lambda_{L}C_{L}N_{0}}h^{(L)}_{a_{L},i,\mu}h^{(L)}_{a_{L},i,\nu}\right)\right]\,.$ (30)
Let $h^{(\ell)}$ be the collection of all the variables $h_{a_{\ell},i,\mu}^{(\ell)}=h_{a_{\ell},i}^{(\ell)}(\mathbf{x}^{\mu})$ where $a_{\ell}=1,\dots,C_{\ell}$, $i=1,\dots,N_{0}$, $\mu=1,\dots,P$. By (4), it is clear that, conditionally on $h^{(\ell-1)}$, $h^{(\ell)}$ is a collection of jointly Gaussian random variables (a Gaussian field) with zero mean and covariance function given by
$\operatorname{Cov}(h_{a,i,\mu},h_{b,j,\nu}|h^{(\ell-1)})=\delta_{a,b}K_{i,j,\mu,\nu}^{(\ell-1)}$
where
$K_{i,j,\mu,\nu}^{(\ell-1)}=\frac{1}{\lambda_{\ell-1}C_{\ell-1}M}\sum_{a_{\ell-1}=1}^{C_{\ell-1}}\sum_{m=-\lfloor M/2\rfloor}^{\lfloor M/2\rfloor}\left(\sum_{r=1}^{N_{0}}T_{m,ir}h^{(\ell-1)}_{a_{\ell-1},r,\mu}\right)\left(\sum_{s=1}^{N_{0}}T_{m,js}h^{(\ell-1)}_{a_{\ell-1},s,\nu}\right).$ (31)
Considering $h^{(\ell)}$ as a vector of dimension $C_{\ell}\times P\times N_{0}$, the conditional distribution of $h^{(\ell)}$ given $h^{(\ell-1)}$ is
$h^{(\ell)}|h^{(\ell-1)}\sim\mathcal{N}(0,\mathbb{1}_{C_{\ell}}\otimes K^{(\ell-1)})$
(see the comment before Lemma 15 for the matrix representation of the tensor $K^{(\ell-1)}$). At this stage, by Gaussian integration, form (30) one gets
$\mathbf{E}[e^{i\bar{\mathbf{s}}_{1:P}^{\top}\mathbf{S}_{1:P}}]$ $=\mathbf{E}\Biggl{[}\det\left((\mathbb{1}_{C_{L}}\otimes K^{(L-1)})^{-1}+\frac{1}{\lambda_{L}C_{L}N_{0}}\mathbb{1}_{C_{L}}\otimes\mathbb{1}_{N_{0}}\otimes(\bar{\mathbf{s}}_{1:P}\bar{\mathbf{s}}_{1:P}^{\top})\right)^{-\frac{1}{2}}$ (32)
$\qquad\qquad\qquad\qquad\det(\mathbb{1}_{C_{L}}\otimes K^{(L-1)})^{-\frac{1}{2}}\Biggr{]}$
$=\mathbf{E}\Biggl{[}\det\left(\mathbb{1}_{N_{0}}\otimes\mathbb{1}_{P}+\frac{1}{\lambda_{L}C_{L}N_{0}}[\mathbb{1}_{N_{0}}\otimes(\bar{\mathbf{s}}_{1:P}\bar{\mathbf{s}}_{1:P}^{\top})]K^{(L-1)}\right)^{-\frac{C_{L}}{2}}\Biggr{]}$
$=\mathbf{E}\Biggl{[}\det\Biggl{(}\mathbb{1}_{N_{0}}+\frac{1}{\lambda_{L}C_{L}N_{0}}\sum_{\mu,\nu}\bar{s}_{\mu}\bar{s}_{\nu}K^{(L-1)}_{\cdot,\mu\nu}\Biggr{)}^{-\frac{C_{L}}{2}}\Biggr{]}\,.$
The last step follows from Lemma 15. Using Eq. (20), we get
$\mathbf{E}[e^{i\bar{\mathbf{s}}_{1:P}^{\top}\mathbf{S}_{1:P}}]$ (33)
$\quad=\mathbf{E}\Biggl{[}\int_{\mathcal{S}^{+}_{N_{0}}}\exp\Biggl{\{}-\frac{1}{2\lambda_{L}N_{0}}\sum_{i,j}Q_{L,ij}\sum_{\mu,\nu}\bar{s}_{\mu}\bar{s}_{\nu}K^{(L-1)}_{ij,\mu\nu}\Biggr{\}}\mathcal{W}_{N_{0}}\Big{(}dQ_{L}|\frac{1}{C_{L}}\mathbb{1}_{N_{0}},C_{L}\Big{)}\Biggr{]}$
$\quad=\mathbf{E}\Biggl{[}\int_{\mathcal{S}^{+}_{N_{0}}}\exp\Biggl{\{}-\frac{1}{2\lambda_{L}\lambda_{L-1}C_{L-1}N_{0}}\sum_{i,j}Q^{*}_{L,ij}\sum_{\mu,\nu}\bar{s}_{\mu}\bar{s}_{\nu}\sum_{a_{L-1}=1}^{C_{L-1}}h^{(L-1)}_{a_{L-1},i,\mu}h^{(L-1)}_{a_{L-1},j,\nu}\Biggr{\}}$
$\qquad\qquad\qquad\qquad\qquad\mathcal{W}_{N_{0}}\Big{(}dQ_{L}|\frac{1}{C_{L}}\mathbb{1}_{N_{0}},C_{L}\Big{)}\Biggr{]}$
where we used Eq. (13).
We could in principle proceed directly integrating out $h^{(L-1)}$ as we did for $h^{(L)}$, Eq. (32); however, in this way we would end up with a Wishart measure for $Q_{L-1}$ dependent on $Q_{L}$. To keep factorized the different contributions, it is easier to change variables with the linear transformation
$\tilde{h}^{(L-1)}_{a_{L-1},i,\mu}=\sum_{j}U^{*}_{L,ij}h^{(L-1)}_{a_{L-1},j,\mu}\,,$ (34)
where $(U^{*}_{L})^{\top}U^{*}_{L}=Q^{*}_{L}$. It is clear that the conditional distribution of $\tilde{h}^{(L-1)}$ given $h^{(L-2)}$ is again Gaussian, more precisely
$\tilde{h}^{(L-1)}|h^{(L-2)}\sim\mathcal{N}(0,\mathbb{1}_{C_{L-1}}\otimes[U^{*}_{L}K^{(L-2)}(U^{*}_{L})^{\top}])\,,$ (35)
with the slight abuse of notation $U^{*}_{L}\equiv U^{*}_{L}\otimes\mathbb{1}_{P}$. Proceeding as before,
$\mathbf{E}[e^{i\bar{\mathbf{s}}_{1:P}^{\top}\mathbf{S}_{1:P}}]$ $=\mathbf{E}\Biggl{[}\int_{(\mathcal{S}^{+}_{N_{0}})^{2}}\exp\Biggl{\{}-\frac{1}{2\lambda_{L}\lambda_{L-1}N_{0}}\operatorname{tr}\Biggl{(}Q_{L-1}\sum_{\mu,\nu}\bar{s}_{\mu}\bar{s}_{\nu}[U^{*}_{L}K^{(L-2)}_{\cdot,\mu\nu}(U^{*}_{L})^{\top}]\Biggr{)}\Biggr{\}}$ (36)
$\qquad\qquad\prod_{\ell\in\{L,L-1\}}\mathcal{W}_{N_{0}}(dQ_{\ell}|\frac{1}{C_{\ell}}\mathbb{1}_{N_{0}},C_{\ell})\Biggr{]}$
$=\mathbf{E}\Biggl{[}\int_{(\mathcal{S}^{+}_{N_{0}})^{2}}\exp\Biggl{\{}-\frac{1}{2\lambda_{L}\lambda_{L-1}\lambda_{L-2}C_{L-2}N_{0}}\sum_{i,j}Q^{*}_{L-1,ij}\sum_{\mu,\nu}\bar{s}_{\mu}\bar{s}_{\nu}$
$\qquad\qquad\sum_{a_{L-2}=1}^{C_{L-2}}h^{(L-2)}_{a_{L-2},i,\mu}h^{(L-2)}_{a_{L-2},j,\nu}\Biggr{\}}\times\prod_{\ell\in\{L,L-1\}}\mathcal{W}_{N_{0}}\Big{(}dQ_{\ell}|\frac{1}{C_{\ell}}\mathbb{1}_{N_{0}},C_{\ell}\Big{)}\Biggr{]}$
where $Q^{*}_{L-1}$ is again given by Eq. (13). Proposition 5 is proven once this procedure is straightforwardly iterated. $\blacklozenge$
Proof of Proposition 7.
The proof of Proposition 7 is identical to the proof of Proposition 6 once one observes that the matrix $\Sigma$, which is $\mathcal{K}_{\mathrm{C}}(Q_{1},\dots,Q_{L})$ for the enlarged dataset $\tilde{X}$, is strictly positive by Lemma 14.